### v2

# Words with factor complexity $2n+1$ and minimal critical exponent

Abstract
Word $\mathbf{G}$} is the fixed point of the morphism $\gamma=[01,2,02]$. In 2019, Shallit and Shur showed that $\mathbf{G}$} has factor complexity $2n+1$. They also showed that $\mathbf{G}$} has critical exponent $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$. They conjectured that this was the least possible critical exponent among words with factor complexity $2n+1$. We confirm their conjecture. The proof, using an intricate case analysis, is by computer. The relevant program generates a ‘human readable’ proof.
Key words: combinatorics on words, factor complexity, critical exponent, computer proof, repetition threshold
1 Introduction
For several decades, the use of computer backtracking has been a standard tool in combinatorics on words. More recently, however, this area has seen a number of trends where computer-assisted proofs come closer to being ‘computer proofs’. One such trend has been seen in the continued success of the Walnut computer package [16, 22], which proves properties of automatic sequences. Over 100 papers using Walnut are listed by Shallit [23]. Of relevance to the current paper, in many situations Walnut can check the critical exponent of the fixed point of a morphism, as in a recent example by Baranwal et. al. [2].
Another instance of the intensive use of computation in proofs has been seen in papers using the ‘template method’ or its variations [1, 9, 10, 15, 21, 5], where one recursively generates ancestors of patterns of interest that could potentially appear in the fixed point of some morphism.
The present paper presents what is, to our knowledge, a new sort of application of computation to combinatorics on words. In analyzing the subject of this paper, a great proliferation of subcases arose, branching to a depth of, for example, subcase A.1.2.2.2 (5 levels of subcases). Some cases involved branching into as many as 7 subcases. Each subcase involved backtrack searches and might have one of several resolutions, possibly including the spawning of new subcases.
Having started with a case analysis by hand, at a certain point we realized that the spawning and resolution of subcases would be best handled by computer. In the end, we wrote a computer program which in turn wrote the desired proof.
The critical exponent of a word is the supremum of the exponents of its factors. In recent years a large number of papers have been written answering questions of this form:
[…]
The various steps in the solution of Dejean’s Conjecture by several authors [12, 18, 17, 6, 3, 11, 20] answered this question for various classes of the form ${\cal L}=\Sigma_{k}^{\omega}$ where $\Sigma_{k}=\{0,1,2,\ldots,k-1\}$.
[…]
In this paper, we find the minimal critical exponent for the class of words with factor complexity $2n+1$. We confirm the 2019 conjecture of Shallit and Shur [24].
2 Word preliminaries
An alphabet is a finite set with elements known as letters. In what follows, we use the alphabet $\Sigma_{3}=\{0,1,2\}$. The set of finite words over $\Sigma_{3}$ is denoted by $\Sigma_{3}^{*}$, and can be regarded as the free semigroup generated by the letters 0, 11, and 22. The length of a word $u$ is the number of occurrences of letters in it and is denoted by $|u|$. Thus, for example, $|01202|=5$. For each $a\in\Sigma_{3}$, we denote the number of occurrences of letter $a$ in $u$ by $|u|_{a}$. Thus, for example, $|01202|_{0}=2,|01202|_{1}=1,|01202|_{2}=2$.
If $u$, $v\in\Sigma_{3}^{*}$, the concatenation $uv$ of $u$ and $v$ consists of the letters of $u$ followed by the letters of $v$. Thus if $u=012$ and $v=02$, we have $uv=01202$. We say that $u$ is a prefix of $uv$ and $v$ is a suffix of $uv$. We also say that $uv$ is a right extension of $u$ by $v$. If $u$, $v$, $w\in\Sigma_{3}^{*}$, we say that $v$ is a factor of uvwuvw. Thus $01$, $120$, and $02$ are respectively a prefix, factor, and suffix of $01202$, and $01202$ is a right extension of $01$ by $202$.
Suppose $u=u_{0}u_{1}u_{2}\cdots u_{n-1}$ where each $u_{i}\in\Sigma_{3}$. We say that positive integer $p$ is a period of $u$ if $u_{i}=u_{i+p}$ whenever $0\leq i\leq i+p\leq n-1$. In this case we say that $u$ is an $r$ power, where $r=n/p$. Thus 3 is a period of word $01201201$, which is a $8/3$ power. Call a word low if it contains no $r$ power for $r\geq 5/2$.
A morphism on $\Sigma_{3}$ is a semigroup homomorphism of $\Sigma_{3}^{*}$. A morphism is determined by its images on letters. Of particular interest is the morphism $\gamma:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ generated by
$\gamma(0)$ $=01$
$\gamma(1)$ $=2$
$\gamma(2)$ $=02$
For conciseness we record a morphism $h$ as $h=[h(0),h(1),h(2)]$. Thus $\gamma=[01,2,02]$. If $|h(a)|>0$ for each $a\in\Sigma_{3}$, we say that $h$ is non-erasing. We use exponentiation to denote iteration of a morphism. Thus $h^{2}(u)=h(h(u))$, $h^{3}(u)=h(h(h(u)))$, etc.
Suppose $u=u_{0}u_{1}u_{2}\cdots u_{n-1}$ where each $u_{i}\in\Sigma_{3}$. The reverse of $u$ is the word $u^{R}=u_{n-1}\cdots u_{2}u_{1}u_{0}$. The reverse of morphism $h=[h(0),h(1),h(2)]$ is defined to be $h^{R}=[(h(0))^{R},(h(1))^{R},(h(2))^{R}]$. Thus, for example, $\gamma^{R}=[10,2,20]$. We see that for a word $u$ and morphism $h$ we have $(h(u))^{R}=h^{R}(u^{R}).$
As well as finite words, we consider right-infinite words over $\Sigma_{3}$. Formally, such a word is a sequence over $\Sigma_{3}$, ${\mathbf{u}}=\{u_{n}\}_{n=0}^{\infty}$ where the $u_{i}\in\Sigma_{3}$. We will use bold-face letters to distinguish right-infinite words from finite words. For each non-negative integer $n$, the length-$n$ prefix of $\mathbf{u}$} is the finite word $u_{0}u_{1}u_{2}\cdots u_{n-1}$. We say that $v\in\Sigma_{3}^{*}$ is a factor of $\mathbf{u}$} if it is a factor of some finite prefix of $\mathbf{u}$}. Call a right-infinite word low if each of its finite prefixes is low.
A final segment of a right-infinite word is the infinite analog of a suffix; a final segment of $\mathbf{u}$} is a right infinite word ${\mathbf{u}}^{\prime}=\{u_{n+i}\}_{n=0}^{\infty}$ for some non-negative integer $i$. We write ${\mathbf{u}}=u{\mathbf{u}}^{\prime}$ where is the prefix of $\mathbf{u}$} of length $i$.
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$. Fix a finite factor $b$ of $\mathbf{w}$}. Word $bu$ is a return word for $b$ in $\mathbf{w}$} if
  • Word bubbub is a factor of $\mathbf{w}$};
  • Whenever $p$ is a prefix of bubbub such that $b$ is a suffix of $p$, then $p=b$ or $p=bub$.
Thus $bu$ is a return word for $b$ in $\mathbf{w}$} if bubbub is a shortest word ‘getting back to’ $b$. If $b$ is a factor of $\mathbf{w}$} then a final segment of $\mathbf{w}$} can be concatenated from the return words of $b$.
The set of all right-infinite words over $\Sigma_{3}$ is denoted by $\Sigma_{3}^{\omega}$. Right infinite words arise in various ways: If $u=u_{0}u_{1}\cdots u_{p-1}$ where $p$ is positive and the $u_{i}$ are letters, then $u^{\infty}=\{u_{i}\}_{i=0}^{\infty}$ where $u_{i}=u_{i\%p}$ where $i\%p$ is the remainder after $i$ is divided by $p$. Thus, for example,
$(012)^{\omega}=012012012012\cdots.$
Another way a right-infinite word arises is via certain morphisms. Let $h$ be a morphism such that $a$ is a prefix of $h(a)$ for some $a\in\Sigma_{3}$. By induction, $h^{n-1}(a)$ is a prefix of $h^{n}(a)$ for each $n$. If $|h^{n}(a)|$ increases without bound, we can define $h^{\omega}(a)$ to be the unique right-infinite word having $h^{n}(a)$ as a prefix for each $n$. We refer to $h^{\omega}(a)$ as a fixed point of $h$.
In this paper we will be particularly interested in the fixed point ${\mathbf{G}}=\gamma^{\omega}(0)$.
A third (non-constructive) way to obtain right-infinite words is via König’s Tree Lemma: A subset $L\subseteq\Sigma_{3}^{*}$ is said to be a factorial language if whenever $\ell$ is in $L$, every factor of $\ell$ is also in $L$. Here is one formulation of König’s Tree Lemma in the combinatorics on words context:
Lemma 1 (König’s Tree Lemma).
Let $L\subseteq\Sigma_{3}^{*}$ be an infinite factorial language. Then there is a right-infinite word ${\mathbf{u}}\in\Sigma_{3}^{\infty}$ such that every prefix of $\mathbf{u}$} is in $L$.
If ${\mathbf{u}}\in\Sigma_{3}^{\omega}$, the factor complexity of $\mathbf{u}$} is the function counting the length $n$ factors of $\mathbf{u}$}. Thus, for example, if ${\mathbf{u}}=(012)^{\omega}$, then the factor complexity of $\mathbf{u}$} is $C(n)$ where
$C(0)$ $=1$
$C(n)$ $=3,n\geq 1.$
Shallit and Shur [24] have shown that $\mathbf{G}$} has factor complexity $2n+1$.
We say that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is eligible if it is low and has factor complexity $2n+1$.
The critical exponent of a finite or right-infinite word is the supremum of those $r$ such that a factor of the word is an $r$ power. Shallit and Shur [24] have shown that $\mathbf{G}$} has critical exponent $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$. They conjecture that this is the least possible critical exponent among words with factor complexity $2n+1$.
We say that right-infinite word ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$, if for any $\epsilon_{1}>0$ there is an $N_{1}$ such that if $u$ is a factor $\mathbf{u}$} of length at least $N_{1}$, then
$\rho_{a}-\epsilon_{1}<\frac{|u|_{a}}{|u|}<\rho_{a}+\epsilon_{1},\text{ for }a\in\Sigma_{3}.$ (1)
Using standard methods, it can be shown that $\mathbf{G}$} has letter frequencies. (For values of the $\rho_{a}$ see the entry for sequence A287104 in the Online Encyclopedia of Integer Sequences; the word $\mathbf{G}$} is a recoding of this sequence.)
3 Main Theorem
Theorem 1 (Main Theorem).
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$. Then the critical exponent of $\mathbf{w}$} is at least $\mu$.
A key ingredient in the proof of this main theorem is a structure theorem.
Theorem 2 (Structure Theorem).
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be low, and let $\mathbf{w}$} have factor complexity $2n+1$. There is a non-erasing morphism $h$ such that either:
  • For every factor $g$ of $\mathbf{G}$}, $h(g)$ is a factor of $\mathbf{w}$}.
  • For every factor $g$ of $\mathbf{G}$}, $h(g^{R})$ is a factor of $\mathbf{w}$}.
The Main Theorem will then follow from this lemma:
Lemma 2.
Let ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ have irrational critical exponent $\nu$, $2<\nu<3.$ Suppose that $\mathbf{u}$} has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$, such that for any $\epsilon_{1}>0$ there is an $N_{1}$ such that if $u$ is a factor $\mathbf{u}$} of length at least $N_{1}$, then
$\rho_{a}-\epsilon_{1}<\frac{|u|_{a}}{|u|}<\rho_{a}+\epsilon_{1},\text{ for }a\in\Sigma_{3}.$ (2)
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ and let $h:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ be a non-erasing morphism. Suppose that for each factor $g$ of $\mathbf{u}$}, either $h(g)$ or $h(g^{R})$ is a factor of $\mathbf{w}$}. Then the critical exponent of $\mathbf{w}$} is at least $\nu$.
Remark 1.
The use of the alphabet $\Sigma_{3}$ and the restriction that $2<\nu<3$ are not essential to this lemma, but simplify the notation in its proof.
Proof of Main Theorem.
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$. Note that $\mu<5/2$. Thus if $\mathbf{w}$} is not low, then the critical exponent of $\mathbf{w}$} is at least $5/2>\mu$, and we are done.
Suppose then that $\mathbf{w}$} is low. By the Structure Theorem, there is a non-erasing morphism $h$ such that for every factor $g$ of $\mathbf{G}$}, either $h(g)$ is a factor of $\mathbf{w}$} or $h(g^{R})$ is a factor of $\mathbf{w}$}. (This is actually weaker than the conclusion of the Structure Theorem).
Word $\mathbf{G}$} has letter frequencies.
It is easy to show that $\mu$ is irrational.
By Lemma 2, then, the critical exponent of ${\mathbf{w}})$ is at least $\mu$. ∎
4 Eligible words and their factors
Lemma 3.
Suppose ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is eligible. Then every final segment of $\mathbf{w}$} contains letters 0, 1, and 2.
Proof.
It suffices to show that every final segment contains letter 0.
Suppose some final segment contains only letters 1 and 2. It must contain factors 12 and 21; otherwise it would end with $1^{\omega}$ or $2^{\omega}$ and therefore contain $5/2$ powers.
It must also contain factor 11 (and, symmetrically, factor 22): If not, assume it starts with 1, replacing it with one of its final segments if necessary. Then it is concatenated from the return words of 1. Since 222 is a $3$ power, the only possible return words of 1 in $\mathbf{w}$} are 12 and 122. The return word 12 can only be used once in this concatenation, or else it appears in the context $2\underline{12}12$, which is a $5/2$ power. Then a final segment is $(122)^{\omega}$, which contains $5/2$ powers. This is a contradiction. Thus 11 must indeed occur.
So far we have shown that every final segment of $\mathbf{w}$} contains factors 11, 12, 21, and 22. Since $\mathbf{w}$} contains $2(2)+1=5$ factors of length 2, there will be exactly one length-2 factor of $\mathbf{w}$} containing a 0. It follows that the only 0 in $\mathbf{w}$} is its first letter. We thus see that $\mathbf{w}$} contains no factor of the form $a0$ where $a\in\Sigma_{3}$.
Consider the following set of 10 length-4 words:
$S=\{1121,1122,1211,1212,1221,2112,2121,2122,2211,2212\}.$
Let $s$ and $t$ be distinct words of $S$. Suppose $u$ is a low word over $\Sigma_{3}$ such that the factor set of $u$ does not include any of 00, 10, 20, $s$, or $t$. For each of the $\binom{10}{2}$ possibilities for $s$ and $t$, a backtrack search shows that the length of $u$ is at most 39. It follows then that $\mathbf{w}$} contains 9 out of 10 words of $S$ as factors. In addition, $\mathbf{w}$} contains a length-4 factor starting with 0. However, $\mathbf{w}$} is eligible, and cannot contain $10$ length-4 factors. This is a contradiction. ∎
Suppose that $\mathbf{w}$} is eligible. Consider the directed graph $D$ with vertex set $\Sigma_{3}$, and with a directed edge $xy$ for $x,y\in\Sigma_{3}$ exactly when $xy$ is a factor of $\mathbf{w}$}. Thus $\mathbf{w}$} can be walked on $D$. Since $\mathbf{w}$} has exactly 5 length-2 factors, $D$ has exactly 5 edges.
By the previous lemma, graph $D$ must be strongly connected. Therefore, it must have at least 3 non-loop edges, and if it has exactly three non-loop edges, they form a directed cycle.
It is convenient to keep track of non-factors of $\mathbf{w}$}: Let $F=\{u\in\Sigma_{3}^{*}:u\text{ is not a factor of }{\mathbf{w}}\}.$ To prove the Structure Theorem, it suffices to consider two cases involving $F$:
  • If $D$ has exactly three non-loop edges (forming a directed cycle) then, permuting $\Sigma_{3}$ if necessary, we may assume that
    $\{01,12,20,00\}\subseteq F.$
  • If $D$ has more than three non-loop edges, permuting the alphabet if necessary,we may assume that
    $\{11,22\}\subseteq F.$
The following lemmas give sufficient conditions for concluding that the Structure Theorem holds:
Lemma 4 (Morphism Lemma).
Suppose that ${\mathbf{w}}\in\{0,1,2\}^{\omega}$ is low. Suppose that $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
  1. 1.
    Word $b$ is a prefix of $a$, which is a prefix of $c$.
  2. 2.
    We have $|b|\geq|c|/2$.
  3. 3.
    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
Then every factor of $h({\mathbf{G}})$ is a factor of $\mathbf{w}$}.
Lemma 5 (Dual Morphism Lemma).
Suppose that ${\mathbf{z}}\in\{0,1,2\}^{\omega}$ is low. Suppose that a final segment of $\mathbf{z}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
  1. 1.
    Word $b$ is a suffix of $a$, which is a suffix of $c$.
  2. 2.
    We have $|b|\geq|c|/2$.
  3. 3.
    Words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
Then if $g$ is any factor of $\mathbf{G}$}, then $h(g^{R})$ is a factor of $\mathbf{z}$}.
5 Proof of the Structure Theorem
We prove the Structure Theorem via an intricate case analysis. Each case is determined by a set $S\subseteq\Sigma_{3}^{*}$ where we assume that $S\subseteq F$, the set of factors omitted by $\mathbf{w}$}. As remarked in the previous section, it suffices to resolve the subcases $S=\{01,12,20,00\}$ and $S=\{11,22\}$.
In a given case, labelled by a set $S$, we proceed in one of three ways. We either:
  1. 1.
    Show that the conditions of the Morphism Lemma or the Dual Morphism Lemma hold, so that the case is resolved;
  2. 2.
    Show that the case cannot arise, so that the case is resolved;
  3. 3.
    Partition the case into subcases.
The proof is thus carried out recursively, in a depth-first way. In situation 1 or 2, a case is resolved. In situation 3, new subcases are added. If at some point all subcases have been resolved (and Hercules beats the hydra) the Structure Theorem has been proved.
The concrete recipe for carrying out 1, 2, and 3 above is as follows:
  1. A.
    If $T\subseteq S$, and case $T$ has been previously resolved, then $S$ is resolved.
  2. B.
    Perform a backtrack search, looking for a low word $w$ of length 250 with no factors in $S$; if no such word exists, the case cannot arise.
  3. C.
    If we find $w$ of length 250, count the factors of $w$ of each length $n$ with $1\leq n\leq 20$, until an $n$ is found such that $w$ has more than $2n+1$ factors of length $n$.
    1. a.
      If no such $n$ is found: We look for a morphism satisfying the conditions of the Morphism Lemma or the Dual Morphism Lemma. To do this, we find factors $b$ of $w$, with $|b|$ up to length 3, such that $b$ has at most 3 recurrent return words in $\mathbf{w}$}. We form a morphism from these 3 return words, and consider it and its conjugates.
      If one of the considered morphisms satisfies the conditions of the Morphism Lemma or the Dual Morphism Lemma the case is resolved.
    2. b.
      If we find $n$ such that $w$ has more than $2n+1$ factors of length $n$, then we classify each such factor $s$ as needed or unneeded. A factor $s$ is needed if a backtrack shows that no low word of length 250 omits all the factors of $S\cup\{s\}$; it is unneeded otherwise.
      1. i.
        If there are more than $2n+1$ needed factors then the condition on the complexity of $\mathbf{w}$} is violated and the case cannot arise.
      2. ii.
        Otherwise, for each unneeded factor $s$ we form a subcase, replacing $S$ by $S\cup\{s\}$.
Remark 2.
The ‘hard-wired’ constants 250 (for backtrack search), 3 (for $|b|$), 20 (depth of the complexity check) are all ad hoc. They were originally taken to be larger, but then tuned.
Proof of the Structure Theorem.
A python implementation of the above scheme in SageMath 9.3 (on a Microsoft Surface with Intel(R) Core(TM) i5-1035G4 CPU at 1.10GHz) resolved the case $S=\{11,22\}$ in about 15 seconds, and the case $S=\{01,12,20,00\}$ in about 4 seconds. ∎
Remark 3.
In fact, the implementation mentioned above kept track of sets $S$ and the reasoning at each of its steps, and produced ‘human readable’ proofs. We give the reasoning that was output for the case $S=\{01,12,20,00\}$ in an appendix. Further, the implementation kept track of all morphisms $h$ used. It may be the case that for some of these morphisms, $h({\mathbf{G}})$ is an eligible word with critical exponent $\mu$.
6 Proofs of Lemmas
Proof of Lemma 2.
The critical exponent of $\mathbf{u}$} is irrational, and therefore is not realized by any of its finite factors. Thus for any $\epsilon_{2}>0$ and any positive integer $N$, word $\mathbf{u}$} has a factor $uuu^{\prime}$ such that $u^{\prime}$ is a prefix of $u$ with $|u^{\prime}|>N$ and
$\frac{|uuu^{\prime}|}{|u|}>\nu-\epsilon_{2}.$ (3)
Since $h$ is non-erasing, we have
$\sum_{a=0}^{2}|h(a)|\rho_{a}\geq\sum_{a=0}^{2}\rho_{a}=1$
and is non-zero. Then
$\lim_{\epsilon\rightarrow 0^{+}}\frac{\sum_{a=0}^{2}|h(a)|(\rho_{a}-\epsilon)}{\sum_{a=0}^{2}|h(a)|(\rho_{a}+\epsilon)}=1$
Given $\epsilon>0$, choose $\epsilon_{1}>0$ such that
$\frac{\sum_{a=0}^{2}|h(a)|(\rho_{a}-\epsilon_{1})}{\sum_{a=0}^{2}|h(a)|(\rho_{a}+\epsilon_{1})}>1-\epsilon.$
Choose $N_{1}$ such that when $|u^{\prime}|>N_{1}$, then
$\rho_{a}-\epsilon_{1}<\frac{|u|_{a}}{|u|}<\rho_{a}+\epsilon_{1},\text{ for }a\in\Sigma_{3}.$
Given $\epsilon_{2}>0$, choose a factor $g=uuu^{\prime}$ of $\mathbf{u}$} with $|u^{\prime}|>N_{1}$ so that (3) holds. (If $h(g^{R})$ is a factor of $\mathbf{w}$}, we replace $g$ by $g^{R}$, redefining $u$ and $u^{\prime}$ accordingly.)
The word $h({\mathbf{u}})$ contains the factor $h(uuu^{\prime})=h(u)h(u)h(u^{\prime})$, and $h(u^{\prime})$ is a prefix of $h(u)$.
Also
$\frac{|h(uuu^{\prime})|}{|h(u)|}$ $=\frac{\sum_{a=0}^{2}|h(a)||uuu^{\prime}|_{a}}{\sum_{a=0}^{2}|h(a)||u|_{a}}$
$\geq\frac{\sum_{a=0}^{2}|h(a)|(\rho_{a}-\epsilon)|uuu^{\prime}|}{\sum_{a=0}^{2}|h(a)|(\rho_{a}+\epsilon)|u|}$
$=\frac{|uuu^{\prime}|}{|u|}\frac{\sum_{a=0}^{2}|h(a)|(\rho_{a}-\epsilon)}{\sum_{a=0}^{2}|h(a)|(\rho_{a}+\epsilon)}$
$>(\nu-\epsilon_{2})(1-\epsilon).$
We see that $\mathbf{w}$} contains factors with exponent arbitrarily close to $\nu$. The result follows. ∎
We use two preliminary lemmas in the proof of the Morphism Lemma (Lemma 4):
Lemma 6.
Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low. Let $a,b,c\in\Sigma_{3}^{+}$, and suppose that a final segment of $\mathbf{w}$} has the form $g({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$, where $g=[a,b,c]$. Suppose further that:
  1. 1.
    Word $b$ is a prefix of $a$, which is a prefix of $c$.
  2. 2.
    We have $|b|\geq|c|/2$.
  3. 3.
    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
Then $\mathbf{u}$} doesn’t contain factors $00$, $11$, 2121, or $22$.
Proof.
We treat the factors one at a time. For each factor we rule out 1-letter right extensions of that factor:
Factor $11$: If $11$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=110$, $v=111$, or $v=112$. In each case, $g(v)$ has the prefix bbbbbb which is a $3$ power. Since $\mathbf{w}$} doesn’t contain $3$ powers, $11$ doesn’t occur in $\mathbf{u}$}.
Factor $00$: If $00$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=000$, $v=001$, or $v=002$. In each case $g(v)$ has the prefix aabaab since $b$ is a prefix of $a$ and $c$. However, aabaab is an $r$ power, $r\geq 5/2$, since $b$ is a prefix of $a$ and $|b|\geq|c|/2\geq|a|/2$.
Factor $21$: If $21$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=211$, $v=210$, or $v=212$. In each case, $g(v)$ has the prefix cbbcbb. This has the suffix sbbsbb, which is an $r$ power, $r\geq 5/2$.
Factor $22$: If $22$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=220$, $v=221$, or $v=222$. In each case, $g(v)$ has the prefix ccbccb, which is an $r$ power, $r\geq 5/2$, since $b$ is a prefix of $c$ and $|b|\geq|c|/2$. ∎
It is convenient to work with $\gamma^{2}=[012,02,0102]$ rather than simply $\gamma$. If $g=[a,b,c]$, then $g\circ\gamma^{2}=[abc,ac,abac]$, and each image of a letter under $g\circ\gamma^{2}$ ends in $c$. This allows us to define $g^{\prime}=c(g\circ\gamma^{2})c^{-1}=[cab,ca,caba]$.
Lemma 7.
Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low. Suppose that $\mathbf{w}$} has the form $g({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $g=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
  1. i.
    Word $b$ is a prefix of $a$, which is a prefix of $c$.
  2. ii.
    We have $|b|\geq|c|/2$.
  3. iii.
    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
Then a final segment of $\mathbf{w}$} has the form $g(\gamma^{2}({\mathbf{v}}))$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$; equivalently, a final segment of $\mathbf{w}$} has the form $g^{\prime}({\mathbf{v}}))$ some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$, where $g^{\prime}=c(g\circ\gamma^{2})c^{-1}=[cab,ca,caba]$. Letting $A=g^{\prime}(0)$, $B=g^{\prime}(1)$, and $C=g^{\prime}(2)$, we have
  1. 1.
    Word $B$ is a prefix of $A$, which is a prefix of $C$.
  2. 2.
    We have $|B|\geq|C|/2$.
  3. 3.
    Words $B$ and $C$ have a common suffix $S$ such that $|S|\geq|B|/2$.
Proof.
Without loss of generality, replacing $\mathbf{w}$} by one of its final segments if necessary, suppose that $\mathbf{w}$} has the form $g({\mathbf{u}})$. By Lemma 6, words $00$ and $11$ are not factors of $\mathbf{u}$}. Therefore, if $2$ is not a factor of some final segment of $\mathbf{u}$}, then that final segment is $(01)^{\omega}$ or $(10)^{\omega}$. This is impossible since $\mathbf{w}$} is low. We conclude that every final segment of $\mathbf{u}$} contains a $2$.
Parse a final segment of $\mathbf{u}$} into blocks starting with $2$, and containing a single $2$. Words $21$ and $22$ are not factors of $\mathbf{u}$}; thus the blocks begin with $20$. Since $00$ and $11$ are not factors, 0’s and $1$’s alternate in the blocks. The block $20101$ is impossible; it would imply a block $z=2010101$ or $z=201012$. In both cases, $g(z)$ begins with cababacababa, since $a$ is a prefix of $c$. However, ababaababa is an $r$ power, $r\geq 5/2$. It follows that a final segment of $\mathbf{u}$} is concatenated from blocks $20$, $201$, and $2010$. Therefore, a final segment of $\mathbf{w}$} is concatenated from blocks $ca=A$, $cab=B$, and $caba=C$, and has the form $g^{\prime}({\mathbf{v}})$ for some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
We see that property (1) certainly holds. Also,
$|C|=|caba|\leq 2|ca|=2|B|$
since $|c|\geq|b|$, establishing (2).
Finally, $C$ and $B$ have the common suffix $S=sa$. However,
$2|s|+|a|\geq|b|+|a|\geq 2|b|\geq|c|.$
Thus
$|S|=|sa|\geq|c|-|s|=|ca|-|sa|=|B|-|S|,$
so that $2|S|\geq|B|$, establishing (3). ∎
Proof of Morphism Lemma.
Define morphisms $g_{n}$ recursively by $g_{1}=g$, and $g_{n+1}=cg_{n}\circ\gamma^{2}c^{-1}$. By induction on the previous lemma, for each positive integer $n$, a final segment of $\mathbf{w}$} has the form $g_{n}({\mathbf{v}})$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
Having a final segment of the form $g_{n+1}({\mathbf{v}})$ is equivalent to having a final segment of the form $g_{n}(\gamma^{2}({\mathbf{v}}))$. By induction on $n$, having a final segment of the form $g_{n}({\mathbf{v}})$ is equivalent to having a final segment of the form $g(\gamma^{2n}({\mathbf{v}}))$. Thus a final segment of $\mathbf{w}$} has the form $g(\gamma^{2n}({\mathbf{v}}))$. ∎
Proof of Dual Morphism Lemma.
Replacing $\mathbf{z}$} by a final segment if necessary, suppose that $\mathbf{z}$} has the form $h({\mathbf{u}})$. Let $S=\{z\in\Sigma_{3}^{*}:h(z)\text{ is a factor of }{\mathbf{z}}\}.$ Let $S^{R}=\{z^{R}:z\in S\}$. Then $S^{R}$ is an infinite factorial language. Using König’s Tree Lemma, choose a word ${\mathbf{z}}^{\prime}\in\Sigma_{3}^{\omega}$ such that every prefix of $\mathbf{z}$$^{\prime}$ is in $S^{R}$.
Define $h^{R}=[h(0)^{R},h(1)^{R},h(2)^{R}]$. Let ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime}).$ Then ${\mathbf{w}}={\mathbf{Z}}$ and $g=h^{R}$ satisfy the conditions of the Morphism Lemma, so that for every $g\in{\mathbf{G}}$, $h^{R}(g)$ is a factor of ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime})$.
Let $p$ be a prefix of $\mathbf{z}$$^{\prime}$ such that $h^{R}(g)$ is a factor of $h^{R}(p).$ Since every prefix of $\mathbf{z}$$^{\prime}$ is in $S^{R}$, write $p=z^{R}$ where $z\in S$. Then $h^{R}(g)$ is a factor of $h^{R}(z^{R})$. Thus $h(g^{R})$ is a factor of $h(z)$. By the definition of $S$, $h(z)$ is a factor of $\mathbf{z}$}, so that $h(g^{R})$ is a factor of $\mathbf{z}$}. ∎
7 A sharpened structure theorem and open problems
Consider the morphism which is the reverse of $\gamma$, namely $\gamma^{R}=[10,2,20]$. Let $\mathbf{G}$$^{\prime}$ be its fixed point
$(\gamma^{R})^{\infty}(2)=20102102010\cdots.$
Then the factors of $\mathbf{G}$$^{\prime}$ are precisely the reverses of the factors of $\mathbf{G}$}. One can easily adjust the proof of the Morphism Lemma to prove
Lemma 8 (Alternate Dual Morphism Lemma).
Suppose that ${\mathbf{z}}\in\{0,1,2\}^{\omega}$ is low. Suppose that a final segment of $\mathbf{z}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
  1. 1.
    Word $b$ is a suffix of $a$, which is a suffix of $c$.
  2. 2.
    We have $|b|\geq|c|/2$.
  3. 3.
    Words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
Then if $g$ is any factor of $\mathbf{G}$$^{\prime}$, then $h(g)$ is a factor of $\mathbf{z}$}.
As remarked earlier, it is possible to keep track of the morphisms used in the proof of the Structure Theorem. Doing so gives the following:
Theorem 3 (Sharpened Structure Theorem).
Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be low, and let $\mathbf{w}$} have factor complexity $2n+1$. One of the following holds:
  • For some non-erasing morphism $h$ in Figure 1, $h(g)$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$};
  • For some non-erasing morphism $h$ in Figure 2, $h(g)$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$$^{\prime}$.
$h(0)$ $h(1)$ $h(2)$
$1210202$ $12102$ $1210202102$
$1210202102$ $1210202$ $121020210202$
$021201212$ $0212012$ $021201212012$
$0201212012$ $0201212$ $020121201212$
$0201212$ $02012$ $0201212012$
$201021010$ $2010210$ $201021010210$
$0102121021$ $0102121$ $010212102121$
$0102121$ $01021$ $0102121021$
$102012020$ $1020120$ $102012020120$
$21021102211021102$ $2102110221102$ $2102110221102110221102$
Figure 1: Potential morphisms $h$ such that some eligible word contains $h(g)$ for each factor $g$ of $\mathbf{G}$}.
$10212$ $0212$ $0210212$
$2120210$ $20210$ $202120210$
$121201210$ $1201210$ $120121201210$
$1201212010$ $1212010$ $121201212010$
$1212010$ $12010$ $1201212010$
$1210120$ $10120$ $101210120$
$202012021$ $2012021$ $201202012021$
$21020$ $1020$ $1021020$
$0210202101$ $0202101$ $020210202101$
$0202101$ $02101$ $0210202101$
$1012102$ $12102$ $121012102$
$101021012$ $1021012$ $102101021012$
$0101202$ $01202$ $0120101202$
$0120101202$ $0101202$ $010120101202$
$010120102$ $0120102$ $012010120102$
$221102210$ $21102210$ $2110221102210$
Figure 2: Potential morphisms $h$ such that some eligible word contains $h(g)$ for each factor $g$ of $\mathbf{G}$$^{\prime}$.
We note in passing that the third morphism in Figure 1, namely
$h=[021201212,0212012,021201212012],$
discovered by the computer implementation of the proof, is a conjugate of
$h_{1}=[201212021,2012021,201212012021].$
Applying the permutation $\sigma=[1,2,0]$ we find
$\sigma(h_{1})=[012020102,0120102,012020120102]=\gamma^{4}.$
Morphism $h$ could thus be replaced in the figure/proof by the identity morphism. However, we did not complicate our proof implementation by searching for simplifications and/or duplications of morphisms.
The sharpened theorem says that the factor set of $\mathbf{w}$} is the same as the factor set of $h({\mathbf{G}})$ for a morphism in Figure 1, or $h({\mathbf{G}}^{\prime})$ for a morphism in Figure 2. It is not known for which, if any, of the $h$ the words $h({\mathbf{G}})$ and $h({\mathbf{G}}^{\prime})$ have factor complexity $2n+1$, and what the critical exponents of these words are. We make the following conjecture:
Conjecture 1.
For each morphism $h$ in Figure 1, the word $h({\mathbf{G}})$ has factor complexity $2n+1$ and critical exponent $\mu$. For each morphism $h$ in Figure 2, the word $h({\mathbf{G}}^{\prime})$ has factor complexity $2n+1$ and critical exponent $\mu$.
Appendix 1: ‘Human readable’ proof of the resolution of the case $S=\{01,12,20,00\}$ of the Structure Theorem
Our Python implementation outputs the following proof:
We assume $F$ includes
$\{01,12,20,00\}$
Every good word longer than 25 must include factor(s)
$\{0221,1021,1022,1102,2102,2110\}$
Author’s comment: In step B of the algorithm, we found a length-250 low word $w$ omitting $\{01,12,20,00\}$. In step C of the algorithm, for $n=4$, we find that $w$ has more than $2n+1$ length-4 factors. Of these, 6 (announced above) are found to be needed factors in step C.b of the algorithm.
Word w must omit a factor from
$\{0210,0211,2210,2211\}$
Author’s comment: These are the 4 unneeded factors found in step C.b of the algorithm. They cannot all be included in $\mathbf{w}$}, because $6+4=10$ would be too many length-4 factors in $\mathbf{w}$}.
This gives rise to 4 cases.
Author’s comment: This is the case division of step C.b.ii.
Case 1: $F$ includes
$\{01,12,20,00,0210\}$
Every good word longer than 68 must include factor(s)
$\{02110,02210,02211,10211,10221,11022,21021,21102,22102,22110\}$
Word w must omit a factor from
$\{11021,21022\}$
This gives rise to 2 cases.
  Case 1.1: $F$ includes
$\{01,12,20,00,0210,11021\}$
Every good word longer than 81 must include factor(s)
$\{02110221022,02110221102,02210211022,02210221102,02211022102,$
10211022102,10211022110,10221021102,10221022110,10221102210,10211022102,10211022110,10221021102,10221022110,10221102210,
11022102110,11022102211,11022110221,21021102210,21021102211,11022102110,11022102211,11022110221,21021102210,21021102211,
21022110221,21102210211,21102210221,21102211022,21022110221,21102210211,21102210221,21102211022,
22102110221,22102211022,22110221021,22110221022}22102110221,22102211022,22110221021,22110221022\}
In this case w must omit factor 02110221021
Author’s comment: Here, because only 1 unneeded factor was found in Step C.b, we don’t make a new subcase.
Author’s comment: It turns out that when we find a length-250 low word $w$ omitting the current list of missing factors, the complexity is $2n+1$ for $n=1,2,\ldots,20$. We therefore (in step C.a) search for a morphism, considering factors $b$ of $w$ of length up to 3, and finding the return words of $b$.
The return words of 210 are among
$\{2102,21021102,210221102,2102110221102\}$
Word 2102 cannot occur twice in a concatenation of these
Author’s comment: If 2102 occurs twice, it occurs in the context $s2102t$, where $s,t$ are two of the return words. However, one checks that each such word $s2102t$ contains an $r$ power, $r\geq 5/2$.
Thus a final segment of w is concatenated from
$\{21021102,210221102,2102110221102\}$
Taking conjugates, a final segment is the image under morphism
$[221102210,21102210,2110221102210]$
which satisfies the Dual Morphism Lemma.
Author’s comment: Since $\mathbf{w}$} has a final segment of the form $h({\mathbf{u}})$, the Dual Morphism Lemma shows that $\mathbf{w}$} has critical exponent at least $\mu$, and the current subcase is resolved.
Case 1.2: $F$ includes
$\{01,12,20,00,0210,21022\}$
The return words of 210 are:
$\{21021102,2102110221102,21021102211021102,2102110221102110221102\}$
Word 21021102 cannot occur twice in a concatenation of these.
    Thus a final segment of w is concatenated from:
$\{2102110221102,21021102211021102,2102110221102110221102\}$
Taking conjugates, a final segment is the image under morphism
[21021102211021102,2102110221102,2102110221102110221102][21021102211021102,2102110221102,2102110221102110221102]
which satisfies the Morphism Lemma.
    Case 2: $F$ includes
$\{01,12,20,00,0211\}$
Every good word longer than 54 must include factor(s)
$\{02102,02210,02211,10210,10221,11021,11022,21021,21022,21102,22102,22110\}$
Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
Author’s comment: We found too many needed factors in step C.b.
Case 3: $F$ includes
$\{01,12,20,00,2210\}$
Every good word longer than 56 must include factor(s)
$\{02102,02110,02211,10210,10211,10221,11021,11022,21021,21022,21102,22110\}$
Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
  Case 4: $F$ includes
$\{01,12,20,00,2211\}$
Every good word longer than 41 must include factor(s)
$\{02102,02110,02210,10210,10211,10221,11021,11022,21021,21022,21102,22102\}$
Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
Author’s comment: The proof in the present case is much shorter than for the resolution of $S=\{11,22\}$. In that longer resolution the computer frequently invokes step A of the algorithm (not used in the present case) avoid repetition of case analysis.
Appendix 2: Python implementation of the case resolution software
The following Python worksheet performs all the calculations mentioned in the paper:
import math
import time
def complexityBreak(w):
# Let the set of factors of w of length n be c[n]. If for some
# 0<i<depth+1 we have |c[i]|>2i+1, return c[i] for the least such
# i. Otherwise return 0.
    depth=20
    for i in range(1,depth+1):
        c=[]
        for j in range(len(w)-i+1):
            if not(w[j:j+i]in c):
                c.append(w[j:j+i])
        if (len(c)>2*i+1):
            c.sort()
            return(c) #
    return(0)
def goodMorphism(L):
# Does L=[a,b,c] satisfy the conditions of the Morphism Lemma?
    def prefix(a,b):
    # Does word b have prefix a?
        if (len(b)<len(a)):
            return(False)
        else:
            return(b[:len(a)]==a)
    def goodsuff(a,c):
    # Do c and a have a common suffix of length |a|/2?
        for i in range(math.ceil(len(a)/2),len(a)+1):
            if (a[-i:]==c[-i:]):
                return(True)
        return(False)
    a=L[0]
    b=L[1]
    c=L[2]
    return(prefix(a,b)and prefix(b,c)and goodsuff(a,c)
           and (2*len(a)>=len(c)))
def dualMorphism(L):
# Does L=[a,b,c] satisfy the conditions of the Dual Morphism Lemma?
    A=L[0][::-1]
    B=L[1][::-1]
    C=L[2][::-1]
    L=[A,B,C]
    return(goodMorphism(L))
def morph(L):
# Returns a conjugate of morphism L which satisfies the Morphism Lemma
# or the Dual Morphism Lemma. If no such conjugate exists, returns 0.
    def morphism_conjugates(L):
    # Given a morphism L (a list of length 3) such that L[0]<=L[1]<=L[2],
    # returns a list of the conjugates of L.
        conjugates=[]
        conjugates.append(L)
        M=L
    # To begin, M is a copy of L. We iteratively conjugate by (prefix)
    # letters
        i=0
        while(i<len(L[0])):
            if ((M[0][0]==M[1][0])and(M[1][0]==M[2][0])):
                # M = [xa,xb,xc], so B=[ax,bx,cx] is a conjugate.
                B=[]
                for j in range(3):
                    B.append(M[j][1:]+M[j][0])
                M=B
                conjugates.append(B)
            i=i+1
    # Again, M is a copy of L. We iteratively conjugate by (suffix)
    # letters
        M=L
        i=1
        while(i<len(L[0])):
            if ((M[0][-1]==M[1][-1])and(M[1][-1]==M[2][-1])):
                # M = [ax,bx,cx], so B=[xa,xb,xc] is a conjugate.
                B=[]
                for j in range(3):
                    B.append(M[j][-1]+M[j][:-1])
                M=B
                conjugates.append(B)
            i=i+1
        return(conjugates)
    C=morphism_conjugates(L)
    for M in C:
        if goodMorphism(M):
            N=[M[1],M[0],M[2]]
            return([1,N])
        if dualMorphism(M):
            N=[M[1],M[0],M[2]]
            return([2,N])
    return([0,[]])
def fhpf(w): # Word w has no suffix of exponent 5/2 or greater
    p=1 #potential period of suffix of exponent 5/2 or greater
    while (5*p<=2*len(w)):
        if (w[-3*p//2:]==w[-5*p//2:-p]):
            return(False)
        p=p+1
    return(True)
def test(w,forbidden_factors):
# This returns true if no suffix of w is in forbidden_factors or has
# exponent at least 5/2
    for f in forbidden_factors:
        if (len(w)>=len(f)):
            if (w[-len(f):]==f):
                return(False)
    return(fhpf(w))
def backtrack(target,forbidden_factors):
# Call a ternary word _low if it contains no factor having exponent
# 5/2 or greater. This routine returns a low word w starting
# with 0 of length target such that w contains no factor in F. If no
# such word exists, it returns the length of the longest low word
# starting with 0 with no factor in forbidden_factors.
    largest_letter=’2’
    w=’0’
    mx=1 # Greatest length attained so far
    while(1==1):
        success=test(w,forbidden_factors)
        if (success):
            if (len(w)>mx):
                mx=len(w)
            if (len(w)==target):
                return(w)
            else:
                w=w+’0’
        else:
            while((len(w)>1) and w[-1]==largest_letter):
                w=w[:-1]
            if (len(w)==1):
#               Search fails. Longest word has length mx.
                return(mx)
            else:
                w=w[:-1]+chr(ord(w[-1])+1)
def good(w,forbidden_factors):
# Call a word _good if it is low and has no factor in
# forbidden_factors.
    for i in range(1,len(w)+1):
        if not(test((w[:i]),forbidden_factors)):
            return(False)
    return(True)
def doubler(n,forbidden_factors):
# This returns the set of all good words of length n. It produces
# the words recursively, testing concatenations of good words of
# lengths ceil(n/2) and floor(n/2).
    if (n==0):
        return([’’])
    if (n==1):
        return([’0’,’1’,’2’])
    G=[]
    for s in doubler(math.ceil(n/2),forbidden_factors):
        for t in doubler(n//2,forbidden_factors):
            u=s+t
            if good(s+t,forbidden_factors):
                G.append(s+t)
    return(G)
def returns(b,depth,forbidden_factors):
# Finds certain return words of b: words bu such that bub is
# good, |u|<=depth, and there are exactly two instances of b in bub,
# namely as prefix and suffix.
    myReturns=[]
    for i in range(depth+1):
        forbidden_factors.append(b) # We seek good words u not
                                    # containing b
        U=doubler(i,forbidden_factors)
        forbidden_factors.pop()
        for u in U:
            if good(b+u+b,forbidden_factors):
                if ((b+u+b)[1:].find(b)==(len(b)+len(u)-1)):
                    myReturns.append(b+u)
    return(myReturns)
def factors(w,n):
# Returns the set of factors of w of length n
    f=[]
    for j in range(len(w)-n+1):
        if not(w[j:j+n] in f):
            f.append(w[j:j+n])
    f.sort()
    return(f)
Morphisms=[]
Dual_Morphisms=[]
longest_b=1
def findMorph(w,k,forbidden_factors):
# Given word w and set of excluded words forbidden_factors, we
# consider factors b # of w of length up to k. Suppose that every
# good word of length v+1 contains b. We find the return words bu.
# If there are exactly three return words (or four return words
# of which the first can only occur once) then any good word has
# a final segment concatenated from the three return words. We
# form a morphism from these return words. We then test whether
# one of its conjugates satisfies the Morphism Lemma or the Dual
# Morphism Lemma,
    global Morphisms
    global Dual_Morphisms
    global longest_b
    n=1
    while(n<=k):
        Fact=factors(w,n)
        for b in Fact:
            forbidden_factors.append(b)
            depth=backtrack(200,forbidden_factors)
            forbidden_factors.pop()
            if(isinstance(depth,int)):
                # Longest good word not containing b has length
                # depth
                Returns=returns(b,depth,forbidden_factors)
                if(len(Returns)>=3):
                    if(len(Returns)<=4):
                            myMorph=[Returns[-3],Returns[-2],Returns[-1]]
                            N=morph(myMorph)
                            if(N[0]==1):
                            # This conjugate of myMorph satisfies
                            # the Morphism Lemma. We check that
                            # Returns[0] cannot appear twice
                            # in a good concatenation of the return
                            # words
                                check=True
                                for p in Returns:
                                    for s in Returns:
                                        if (good(p+Returns[0]+s,
                                                 forbidden_factors)):
                                            check=False
                                if (check):
                                    print(indent+’The return words ’+
                                          ’of ’,b,’ are:’+’\\’+’\\’)
                                    print(Returns,’\\’+’\\’)
                                    print(indent+’Word ’,Returns[0],
                                    ’cannot occur twice in a ’+
                                    ’concatenation of these.\\’+’\\’)
                                    print(indent+’Thus a final ’+
                                    ’segment of w is concatenated ’+
                                    ’from:’,’\\’+’\\’)
                                    print(myMorph,’\\’+’\\’)
                                    print(indent+’Taking conjugates,’+
                                    ’ a final segment is the image ’+
                                    ’under morphism’,’\\’+’\\’)
                                    print(N[1],’\\’+’\\’)
                                    if (not(N[1]in Morphisms)):
                                        Morphisms.append(N[1])
                                    print(indent+’which satisfies’+
                                    ’ the Morphism Lemma.\\’+’\\’)
                                    return()
                            if(N[0]==2): # This conjugate of myMorph
                                         # satisfies the Dual Morphism Lemma
                                check=True
                                for p in Returns:
                                    for s in Returns:
                                        if (good(p+Returns[0]+s,
                                                 forbidden_factors)):
                                            check=False
                                if (check):
                                    print(indent+’The return words’+
                                    ’ of ’,b,’ are among ’,
                                    Returns,’\\’+’\\’)
                                    print(indent+’Word ’,Returns[0],
                                    ’cannot occur twice in a ’+
                                    ’concatenation of these’,’\\’+
                                    print(indent+’Thus a final ’
                                    +’segment of w is ’+
                                    ’ concatenated from ’,myMorph,
                                    print(indent+’Taking conjugates,’+
                                    ’ a final segment is the image ’+
                                    ’under morphism’,’\\’+’\\’)
                                    print(N[1],’\\’+’\\’)
                                    if (not(N[1]in Dual_Morphisms)):
                                        Dual_Morphisms.append(N[1])
                                    print(indent+’which satisfies’+
                                    ’ the Dual Morphism ’+
                                    ’Lemma.’,’\\’+’\\’)
                                    return()
        n=n+1
        longest_b=max(longest_b,n)
    print(indent+’No morphism found up to length ’, k,’\\’+’\\’)
    return()
indent = ’’
def resolveCase(forbidden_factors,caseString):
    global indent
    global resolved_cases
    global resolved_case_labels
    indent=’~’
    for i in range(len(caseString)):
        indent=indent+’~’
    k=3 # Maximum length of b considered in return words bu
    if (caseString==’’):
        print(indent+’We assume F includes ’,forbidden_factors,
    w=backtrack(250,forbidden_factors)
    if (isinstance(w,int)):
        print(’No good word longer than ’,w,’ avoids these factors.’
        F=[]
        for f in forbidden_factors:
            F.append(f)
        resolved_cases.append(F)
        resolved_case_labels.append(caseString)
        print(resolved_cases)
        print(resolved_case_labels)
        return()
#   The case being resolved is labelled by forbidden_factors.
#   However, we may find that additional factors are also necessarily
#   avoided in this case. We will sharpen our further backtracking
# searches within this case by avoiding these additional factors.
#   To put it another way, when there is exactly one unneeded factor,
#   we would get one new subcase. Since there is no branching, it is
#   more intuitive to simply say the subcase is the "same" case, but
#   we specify additional avoided factors in factorsToAvoid. The "base
#   case" is still labelled by the set forbidden_factors
    factorsToAvoid=[]
    for f in forbidden_factors:
        factorsToAvoid.append(f)
    unneeded_factors=[]
    while(len(unneeded_factors)<2):
        w=backtrack(250,factorsToAvoid)
        N=complexityBreak(w)
        if (N==0): #complexity is good
            M=findMorph(w,k,factorsToAvoid)
            F=[]
            for f in forbidden_factors:
                F.append(f)
            resolved_cases.append(F)
            resolved_case_labels.append(caseString)
            return()
        else:
            Mx=0
            needed_factors=[]
            unneeded_factors=[]
            for s in N:
                factorsToAvoid.append(s)
                w=backtrack(250,factorsToAvoid)
                factorsToAvoid.pop()
                if (isinstance(w,int)):
                    Mx=max(Mx,w)
                    needed_factors.append(s)
                else:
                    unneeded_factors.append(s)
            if(len(needed_factors)>0):
                print(indent+’Every good word longer than ’,Mx,
                ’must include factor(s)’,’\\’+’\\’)
                print(needed_factors,’\\’+’\\’)
                if(len(needed_factors)>2*len(N[0])+1):
                    print(indent+’Word w has ’,len(needed_factors),
                    ’> 2 x’,len(N[0]),’+ 1 factors of length ’,
                    len(N[0]) ,’. This is impossible.’,’\\’+’\\’)
                    F=[]
                    for f in forbidden_factors:
                        F.append(f)
                    resolved_cases.append(F)
                    resolved_case_labels.append(caseString)
                    return()
            if (len(unneeded_factors)==1):
                print(indent+’In this case w must omit factor ’,
                unneeded_factors[0],
                s=unneeded_factors[0]
                factorsToAvoid.append(s)
    unneeded_factors.sort()
    j=0
    print(indent+’Word w must omit a factor from’,’\\’+’\\’)
    print(unneeded_factors,’\\’+’\\’)
    print(indent+’This gives rise to ’,len(unneeded_factors),
    ’ cases.’,’\\’+’\\’)
    for s in unneeded_factors:
        if (caseString==’’):
            subcase=chr(ord(’1’)+j)
        else:
            subcase=caseString+’.’+chr(j+49)
        j=j+1
        factorsToAvoid.append(s)
        print(indent+’Case ’+subcase+’: F includes ’,factorsToAvoid,
        # Test whether case was resolved previously
        resolved=False
        for c in resolved_cases:
            resolved=True
            for d in c:
                if not(d in factorsToAvoid):
                    resolved=False
                    break
            if (resolved):
                i=resolved_cases.index(c)
                print(’This was previously resolved in Case ’+
                resolved_case_labels[i]+’.’,’\\’+’\\’)
                break
        if(not(resolved)):
            resolveCase(factorsToAvoid,subcase)
            factorsToAvoid.pop()
    F=[]
    for f in forbidden_factors:
        F.append(f)
    resolved_cases.append(F)
    resolved_case_labels.append(caseString)
    return()
print(’Here is the resolution of S={22,11}:’)
print(’ ’)
resolved_cases=[]
resolved_case_labels=[]
startTime=time.time()
resolveCase([’22’,’11’],’’)
endTime=time.time()
print(’Computation took ’,(endTime-startTime),’ seconds.’)
print(’ ’)
print(’ ’)
print(’ ’)
print(’Here is the resolution of S={01,12,20,00}:’)
print(’ ’)
resolved_cases=[]
resolved_case_labels=[]
startTime=time.time()
resolveCase([’01’,’12’,’20’,’$0$’],’’)
endTime=time.time()
print(’Computation took ’,(endTime-startTime),’ seconds.’)
print(’ ’)
print(’ ’)
print(’ ’)
print(’Here are the morphisms satisfying the Morphism Lemma used ’+
’in the case resolutions:’)
print(’ ’)
for m in Morphisms:
    print(m)
print(’ ’)
print(’ ’)
print(’ ’)
print(’Here are the morphisms satisfying the Dual Morphism Lemma ’+
’used in the case resolutions:’)
print(’ ’)
for m in Dual_Morphisms:
    print(m)
print(’ ’)
print(’ ’)
print(’ ’)
print(’Here is the length of the longest word in the backtracks of’+
’ Lemma 2:’)
S=[’1121’, ’1122’, ’1211’, ’1212’, ’1221’, ’2112’,
’2121’, ’2122’, ’2211’, ’2212’]
F=[’10’,’20’,’$0$’]
longest=0
for s in S:
        for t in S:
            if (s!=t):
                F.append(s)
                F.append(t)
                w=backtrack(500,F)
                longest=max(w,longest)
                F.pop()
                F.pop()
print(’ ’)
print(longest)

### v3

@@
 Abstract
-Word $\mathbf{G}$} is the fixed point of the morphism $\gamma=[01,2,02]$.
+Let the word $\mathbf{G}$} be the fixed point of the morphism $\gamma$ sending 0 to $01$, 11 to $2$, and $2$ to $02$.
 In 2019, Shallit and Shur showed that $\mathbf{G}$} has factor complexity $2n+1$.
 They also showed that $\mathbf{G}$} has critical exponent $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$.
-They conjectured that this was the least possible critical exponent among words with factor complexity $2n+1$.
+They conjectured that this was the least possible critical exponent among the words with factor complexity $2n+1$.
 We confirm their conjecture.
-The proof, using an intricate case analysis, is by computer.
-The relevant program generates a ‘human readable’ proof.
+We anticipate that our method, including an intricate case analysis by computer, will have wider application.
 Key words: combinatorics on words, factor complexity, critical exponent, computer proof, repetition threshold
 1 Introduction
+The critical exponent of a word is the supremum of the exponents of its factors.
+(Definitions of basic terms in combinatorics on words, such as factors, exponents, etc., follow in the next section.) In recent years a large number of papers have been written answering questions of the following form:
+[…]
+[…]
+For example, the minimum critical exponent for a Sturmian word is
+$2+\varphi=3.618033988749\cdots$
+where $\varphi$ is the golden ratio.
+This is realized by the Fibonacci word, the fixed point of the morphism $\Phi$ sending 0 to $01$ and 11 to 0.
+One starting point for the study of Sturmian words and their generalizations is the relevant chapter in Allouche and Shallit’s book [1].
+Words of factor complexity $2n+1$ are considered in the paper of Cassaigne, Labbé, and Leroy [6].
+Such words are connected to the well-studied Arnoux-Rauzy words: the Arnoux-Rauzy words over a $k$-letter alphabet have factor complexity (k1)n+1(k-1)n+1.
+Note that the Arnoux-Rauzy words over a binary alphabet are just the Sturmian words.
+In the present paper, we find the minimal critical exponent for the class of words with factor complexity $2n+1$.
+This confirms the 2019 conjecture of Shallit and Shur [28].
+Taken together with Carpi and De Luca’s result [5] for Sturmian words, this gives the beginnings of a Dejean-type theorem for Arnoux-Rauzy words: we know the minimal critical exponent in the cases $k=2$ and $k=3$.
+To find the minimal critical exponent for binary rich words, it was necessary to prove the following structure theorem:
+Theorem 1.
+[9] Let ${\mathbf{w}}\in\Sigma_{2}^{\omega}$ be a $14/5$-power-free rich word.
+For every $n\geq 1$, a suffix of $\mathbf{w}$} has the form $f(\phi^{n}(w_{n}))$ or $f(g(\phi^{n}(w_{n})))$ for some word $w_{n}\in\Sigma_{3}^{\omega}$ where
+$f(0)=$ $0$ $g(0)=$ $011$ $\phi(0)=$ $01$
+$f(1)=$ $01$ $g(1)=$ $0121$ $\phi(1)=$ $02$
+$f(2)=$ $011$ $g(2)=$ $012121$ $\phi(2)=$ $022.$
+Paraphrasing, if we let $\mathbf{H}$} be the fixed point of $\phi$, then the infinite $14/5$-power-free rich binary words look like one of $h_{1}({\mathbf{H}})$ and $h_{2}({\mathbf{H}})$, where $h_{1}=f$, and $h_{2}=f\circ g$.
+In the present paper, we prove the analogous structure theorem for the infinite words with factor complexity $2n+1$, and not containing $r$-powers for $r\geq 5/2$.
+However, instead of the two possible morphisms $h_{1}$ and $h_{2}$ in the above-mentioned theorem, there are several possible morphisms.
+(See Figures 1 and 2 and following.) The case analysis is therefore complex.
+(Interestingly, Shallit and Shur [28] prove that the minimum critical exponent for the infinite words with factor complexity $2n$ is exactly $5/2$.)
 For several decades, the use of computer backtracking has been a standard tool in combinatorics on words.
@@
 Over 100 papers using Walnut are listed by Shallit [27].
-Of relevance to the current paper, in many situations Walnut can check the critical exponent of the fixed point of a morphism, as in a recent example by Baranwal et.
-al.
+Of relevance to the current paper, in many situations Walnut can check the critical exponent of the fixed point of a morphism, as in a recent example by Baranwal et al.
 [3].
@@
 In the end, we wrote a computer program which in turn wrote the desired proof.
-The critical exponent of a word is the supremum of the exponents of its factors. In recent years a large number of papers have been written answering questions of this form:
-[…]
-The various steps in the solution of Dejean’s Conjecture by several authors [12, 18, 17, 6, 3, 11, 20] answered this question for various classes of the form ${\cal L}=\Sigma_{k}^{\omega}$ where $\Sigma_{k}=\{0,1,2,\ldots,k-1\}$.
-[…]
-In this paper, we find the minimal critical exponent for the class of words with factor complexity $2n+1$. We confirm the 2019 conjecture of Shallit and Shur [24].
 2 Word preliminaries
@@
 If $u$, $v$, $w\in\Sigma_{3}^{*}$, we say that $v$ is a factor of uvwuvw.
+If word $y$ is not a factor of word $z$ we say that $z$ omits the factor $y$.
 Thus $01$, $120$, and $02$ are respectively a prefix, factor, and suffix of $01202$, and $01202$ is a right extension of $01$ by $202$.
+Word $01202$ omits the factor $10$.
 Suppose $u=u_{0}u_{1}u_{2}\cdots u_{n-1}$ where each $u_{i}\in\Sigma_{3}$.
@@
 In this case we say that $u$ is an $r$ power, where $r=n/p$.
-Thus 3 is a period of word $01201201$, which is a $8/3$ power.
+Thus 3 is a period of the word $01201201$, which is an $8/3$ power.
 Call a word low if it contains no $r$ power for $r\geq 5/2$.
@@
 The reverse of $u$ is the word $u^{R}=u_{n-1}\cdots u_{2}u_{1}u_{0}$.
-The reverse of morphism $h=[h(0),h(1),h(2)]$ is defined to be $h^{R}=[(h(0))^{R},(h(1))^{R},(h(2))^{R}]$.
+The reverse of the morphism $h=[h(0),h(1),h(2)]$ is defined to be $h^{R}=[(h(0))^{R},(h(1))^{R},(h(2))^{R}]$.
 Thus, for example, $\gamma^{R}=[10,2,20]$.
-We see that for a word $u$ and morphism $h$ we have $(h(u))^{R}=h^{R}(u^{R}).$
-As well as finite words, we consider right-infinite words over $\Sigma_{3}$.
-Formally, such a word is a sequence over $\Sigma_{3}$, ${\mathbf{u}}=\{u_{n}\}_{n=0}^{\infty}$ where the $u_{i}\in\Sigma_{3}$.
-We will use bold-face letters to distinguish right-infinite words from finite words.
+We see that for a word $u$ and the morphism $h$ we have $(h(u))^{R}=h^{R}(u^{R}).$
+As well as the finite words, we consider the right-infinite words over $\Sigma_{3}$.
+Formally, such a word is a sequence ${\mathbf{u}}=\{u_{n}\}_{n=0}^{\infty}$ where each $u_{i}$ is a single letter.
+We use bold-face letters to distinguish right-infinite words from finite words.
 For each non-negative integer $n$, the length-$n$ prefix of $\mathbf{u}$} is the finite word $u_{0}u_{1}u_{2}\cdots u_{n-1}$.
@@
 Call a right-infinite word low if each of its finite prefixes is low.
-A final segment of a right-infinite word is the infinite analog of a suffix; a final segment of $\mathbf{u}$} is a right infinite word ${\mathbf{u}}^{\prime}=\{u_{n+i}\}_{n=0}^{\infty}$ for some non-negative integer $i$.
-We write ${\mathbf{u}}=u{\mathbf{u}}^{\prime}$ where is the prefix of $\mathbf{u}$} of length $i$.
-Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$. Fix a finite factor $b$ of $\mathbf{w}$}. Word $bu$ is a return word for $b$ in $\mathbf{w}$} if
-  • Word bubbub is a factor of $\mathbf{w}$};
+We extend the notaion of suffix to right-infinite words; a suffix of $\mathbf{u}$} is a right-infinite word ${\mathbf{u}}^{\prime}=\{u_{n+i}\}_{n=0}^{\infty}$ for some non-negative integer $i$.
+We write ${\mathbf{u}}=u{\mathbf{u}}^{\prime}$ where $u$ is the prefix of $\mathbf{u}$} of length $i$.
+Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$. Fix a finite factor $b$ of $\mathbf{w}$}. The word $bu$ is a $b$-block of $\mathbf{w}$} if
+  • The word bubbub is a factor of $\mathbf{w}$};
   • Whenever $p$ is a prefix of bubbub such that $b$ is a suffix of $p$, then $p=b$ or $p=bub$.
-Thus $bu$ is a return word for $b$ in $\mathbf{w}$} if bubbub is a shortest word ‘getting back to’ $b$.
-If $b$ is a factor of $\mathbf{w}$} then a final segment of $\mathbf{w}$} can be concatenated from the return words of $b$.
+If $b$ appears infinitely often as a factor of $\mathbf{w}$} then a suffix of $\mathbf{w}$} can be concatenated from the $b$-blocks of $\mathbf{w}$}.
 The set of all right-infinite words over $\Sigma_{3}$ is denoted by $\Sigma_{3}^{\omega}$.
-Right infinite words arise in various ways: If $u=u_{0}u_{1}\cdots u_{p-1}$ where $p$ is positive and the $u_{i}$ are letters, then $u^{\infty}=\{u_{i}\}_{i=0}^{\infty}$ where $u_{i}=u_{i\%p}$ where $i\%p$ is the remainder after $i$ is divided by $p$.
+Right-infinite words arise in various ways: If $u=u_{0}u_{1}\cdots u_{p-1}$ where $p$ is positive and the $u_{i}$ are letters, then $u^{\infty}=\{u_{i}\}_{i=0}^{\infty}$ where $u_{i}=u_{i\bmod p}$.
 Thus, for example,
@@
 If $|h^{n}(a)|$ increases without bound, we can define $h^{\omega}(a)$ to be the unique right-infinite word having $h^{n}(a)$ as a prefix for each $n$.
-We refer to $h^{\omega}(a)$ as a fixed point of $h$.
-In this paper we will be particularly interested in the fixed point ${\mathbf{G}}=\gamma^{\omega}(0)$.
-A third (non-constructive) way to obtain right-infinite words is via König’s Tree Lemma: A subset $L\subseteq\Sigma_{3}^{*}$ is said to be a factorial language if whenever $\ell$ is in $L$, every factor of $\ell$ is also in $L$.
-Here is one formulation of König’s Tree Lemma in the combinatorics on words context:
-Lemma 1 (König’s Tree Lemma).
+For a right-infinite word $\mathbf{u}$} over $\Sigma_{3}$, we define $h({\mathbf{u}})$ to be the right-infinite word having $h(u)$ as a prefix for each prefix $u$ of $\mathbf{u}$}; thus $h^{\omega}(a)$ is a fixed point of $h$.
+In this paper we are particularly interested in the fixed point ${\mathbf{G}}=\gamma^{\omega}(0)$.
+A third (non-constructive) way to obtain right-infinite words is via König’s lemma [18].
+Here is one formulation of König’s lemma in the combinatorics on words context: A subset $L\subseteq\Sigma_{3}^{*}$ is said to be a factorial language if whenever $\ell$ is in $L$, every factor of $\ell$ is also in $L$.
+Lemma 1 (König’s lemma).
 Let $L\subseteq\Sigma_{3}^{*}$ be an infinite factorial language. Then there is a right-infinite word ${\mathbf{u}}\in\Sigma_{3}^{\infty}$ such that every prefix of $\mathbf{u}$} is in $L$.
@@
 They conjecture that this is the least possible critical exponent among words with factor complexity $2n+1$.
-We say that right-infinite word ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$, if for any $\epsilon_{1}>0$ there is an $N_{1}$ such that if $u$ is a factor $\mathbf{u}$} of length at least $N_{1}$, then
+We say that the right-infinite word ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$, if for all $\epsilon_{1}>0$ there is an $N_{1}$ such that if $u$ is a factor $\mathbf{u}$} of length at least $N_{1}$, then
 $\rho_{a}-\epsilon_{1}<\frac{|u|_{a}}{|u|}<\rho_{a}+\epsilon_{1},\text{ for }a\in\Sigma_{3}.$ (1)
 Using standard methods, it can be shown that $\mathbf{G}$} has letter frequencies.
-(For values of the $\rho_{a}$ see the entry for sequence A287104 in the Online Encyclopedia of Integer Sequences; the word $\mathbf{G}$} is a recoding of this sequence.)
+(For values of the $\rho_{a}$ see the entry for sequence A287104 in the On-Line Encyclopedia of Integer Sequences; the word $\mathbf{G}$} is a recoding of this sequence.)
 3 Main Theorem
 Theorem 2 (Main Theorem).
-Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$. Then the critical exponent of $\mathbf{w}$} is at least $\mu$.
+Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$.
+Then the critical exponent of $\mathbf{w}$} is at least $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$.
 A key ingredient in the proof of this main theorem is a structure theorem.
 Theorem 3 (Structure Theorem).
-Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be low, and let $\mathbf{w}$} have factor complexity $2n+1$. There is a non-erasing morphism $h$ such that either:
-  • For every factor $g$ of $\mathbf{G}$}, $h(g)$ is a factor of $\mathbf{w}$}.
-  • For every factor $g$ of $\mathbf{G}$}, $h(g^{R})$ is a factor of $\mathbf{w}$}.
-The Main Theorem will then follow from this lemma:
+Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be eligible. There is a non-erasing morphism $h$ such that either
+  • For every factor $g$ of $\mathbf{G}$}, the word $h(g)$ is a factor of $\mathbf{w}$}.
+  • For every factor $g$ of $\mathbf{G}$}, the word $h(g^{R})$ is a factor of $\mathbf{w}$}.
+The main theorem follows from this lemma:
 Lemma 2.
-Let ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ have irrational critical exponent $\nu$, $2<\nu<3.$ Suppose that $\mathbf{u}$} has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$, such that for any $\epsilon_{1}>0$ there is an $N_{1}$ such that if $u$ is a factor $\mathbf{u}$} of length at least $N_{1}$, then
-$\rho_{a}-\epsilon_{1}<\frac{|u|_{a}}{|u|}<\rho_{a}+\epsilon_{1},\text{ for }a\in\Sigma_{3}.$ (2)
+Let ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ have irrational critical exponent $\nu$, with $2<\nu<3.$ Suppose that $\mathbf{u}$} has letter frequencies $\rho_{0}$, $\rho_{1}$, $\rho_{2}$.
 Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ and let $h:\Sigma_{3}^{*}\rightarrow\Sigma_{3}^{*}$ be a non-erasing morphism.
@@
 Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ have factor complexity $2n+1$.
-Note that $\mu<5/2$.
-Thus if $\mathbf{w}$} is not low, then the critical exponent of $\mathbf{w}$} is at least $5/2>\mu$, and we are done.
+Recall that $\mu=2+\frac{1}{\lambda^{2}-1}=2.4808726\cdots$, where $\lambda=1.7548777$ is the real zero of $x^{3}-2x+x-1=0$.
+Thus $\mu<5/2$.
+If $\mathbf{w}$} is not low, then the critical exponent of $\mathbf{w}$} is at least $5/2>\mu$, and we are done.
 Suppose then that $\mathbf{w}$} is low.
@@
 (This is actually weaker than the conclusion of the structure theorem).
-Word $\mathbf{G}$} has letter frequencies.
+The word $\mathbf{G}$} has letter frequencies.
 It is easy to show that $\mu$ is irrational.
-By Lemma 2, then, the critical exponent of ${\mathbf{w}})$ is at least $\mu$. ∎
+By Lemma 2, then, the critical exponent of $\mathbf{w}$} is at least $\mu$.
+∎
 4 Eligible words and their factors
+We defer the proof of the following lemma.
 Lemma 3.
-Suppose ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is eligible. Then every final segment of $\mathbf{w}$} contains letters 0, 1, and 2.
-Proof.
-It suffices to show that every final segment contains letter 0.
-Suppose some final segment contains only letters 1 and 2. It must contain factors 12 and 21; otherwise it would end with $1^{\omega}$ or $2^{\omega}$ and therefore contain $5/2$ powers.
-It must also contain factor 11 (and, symmetrically, factor 22): If not, assume it starts with 1, replacing it with one of its final segments if necessary.
-Then it is concatenated from the return words of 1.
-Since 222 is a $3$ power, the only possible return words of 1 in $\mathbf{w}$} are 12 and 122.
-The return word 12 can only be used once in this concatenation, or else it appears in the context $2\underline{12}12$, which is a $5/2$ power.
-Then a final segment is $(122)^{\omega}$, which contains $5/2$ powers.
-This is a contradiction.
-Thus 11 must indeed occur.
-So far we have shown that every final segment of $\mathbf{w}$} contains factors 11, 12, 21, and 22.
-Since $\mathbf{w}$} contains $2(2)+1=5$ factors of length 2, there will be exactly one length-2 factor of $\mathbf{w}$} containing a 0.
-It follows that the only 0 in $\mathbf{w}$} is its first letter.
-We thus see that $\mathbf{w}$} contains no factor of the form $a0$ where $a\in\Sigma_{3}$.
-Consider the following set of 10 length-4 words:
-$S=\{1121,1122,1211,1212,1221,2112,2121,2122,2211,2212\}.$
-Let $s$ and $t$ be distinct words of $S$.
-Suppose $u$ is a low word over $\Sigma_{3}$ such that the factor set of $u$ does not include any of 00, 10, 20, $s$, or $t$.
-For each of the $\binom{10}{2}$ possibilities for $s$ and $t$, a backtrack search shows that the length of $u$ is at most 39.
-It follows then that $\mathbf{w}$} contains 9 out of 10 words of $S$ as factors.
-In addition, $\mathbf{w}$} contains a length-4 factor starting with 0.
-However, $\mathbf{w}$} is eligible, and cannot contain $10$ length-4 factors.
-This is a contradiction.
-∎
+Suppose ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is eligible. Then every suffix of $\mathbf{w}$} contains letters 0, 1, and 2.
 Suppose that $\mathbf{w}$} is eligible.
 Consider the directed graph $D$ with vertex set $\Sigma_{3}$, and with a directed edge $xy$ for $x,y\in\Sigma_{3}$ exactly when $xy$ is a factor of $\mathbf{w}$}.
-Thus $\mathbf{w}$} can be walked on $D$.
-Since $\mathbf{w}$} has exactly 5 length-2 factors, $D$ has exactly 5 edges.
-By the previous lemma, graph $D$ must be strongly connected. Therefore, it must have at least 3 non-loop edges, and if it has exactly three non-loop edges, they form a directed cycle.
-It is convenient to keep track of non-factors of $\mathbf{w}$}: Let $F=\{u\in\Sigma_{3}^{*}:u\text{ is not a factor of }{\mathbf{w}}\}.$ To prove the Structure Theorem, it suffices to consider two cases involving $F$:
+Then there is a directed path on $D$ labeled with the letters of $\mathbf{w}$}.
+Since $\mathbf{w}$} has exactly 5 length-2 factors, the graph $D$ has exactly 5 edges.
+By the previous lemma, the graph $D$ must be strongly connected; that is, there is a directed path between any two vertices of $D$.
+Therefore, $D$ must have at least 3 non-loop edges, and if it has exactly three non-loop edges, they form a directed cycle.
+It is convenient to keep track of words which are not factors of $\mathbf{w}$}: Let $F=\{u\in\Sigma_{3}^{*}:u\text{ is not a factor of }{\mathbf{w}}\}.$ To prove the structure theorem, it suffices to consider two cases involving $F$.
   • If $D$ has exactly three non-loop edges (forming a directed cycle) then, permuting $\Sigma_{3}$ if necessary, we may assume that
@@
     $\{11,22\}\subseteq F.$
-The following lemmas give sufficient conditions for concluding that the Structure Theorem holds:
+The following technical lemmas give sufficient conditions for concluding that the structure theorem holds:
 Lemma 4 (Morphism Lemma).
-Suppose that ${\mathbf{w}}\in\{0,1,2\}^{\omega}$ is low.
-Suppose that $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
+Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low.
+Suppose that some suffix of $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that
   1. 1.
-    Word $b$ is a prefix of $a$, which is a prefix of $c$.
+    The word $b$ is a prefix of $a$, which is a prefix of $c$.
   2. 2.
@@
   3. 3.
-    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
-Then every factor of $h({\mathbf{G}})$ is a factor of $\mathbf{w}$}.
+    The words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
+If $g$ is any factor of $\mathbf{G}$}, then $h(g)$ is a factor of $\mathbf{w}$}.
 Lemma 5 (Dual Morphism Lemma).
-Suppose that ${\mathbf{z}}\in\{0,1,2\}^{\omega}$ is low.
-Suppose that a final segment of $\mathbf{z}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
+Suppose that ${\mathbf{z}}\in\Sigma_{3}^{\omega}$ is low.
+Suppose that a suffix of $\mathbf{z}$} has the form $h({\mathbf{u}})$ for some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$, where $h=[a,b,c]$ for some $a,b,c\in\Sigma_{3}^{+}$ such that
   1. 1.
-    Word $b$ is a suffix of $a$, which is a suffix of $c$.
+    The word $b$ is a suffix of $a$, which is a suffix of $c$.
   2. 2.
@@
   3. 3.
-    Words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
-Then if $g$ is any factor of $\mathbf{G}$}, then $h(g^{R})$ is a factor of $\mathbf{z}$}.
+    The words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
+If $g$ is any factor of $\mathbf{G}$}, then $h(g^{R})$ is a factor of $\mathbf{z}$}.
+Lemma 6 (First Parsing Lemma).
+Let $S\subseteq\Sigma_{3}^{*}$ and let $b\in\Sigma_{3}^{*}$.
+Let $d$ be a non-negative integer.
+Suppose that every low word of length $d+1$ contains a factor from $S\cup\{b\}$.
+Then if $\mathbf{w}$} is an eligible right-infinite word which omits the words of $S$, a suffix of $\mathbf{w}$} can be concatenated from $b$-blocks of the form $bu$, where $|u|\leq d$.
+Lemma 7 (Second Parsing Lemma).
+Let $S,B\subseteq\Sigma_{3}^{*}$.
+Let $\mathbf{w}$} be an eligible right-infinite word which omits the words of $S$, such that $\mathbf{w}$} is concatenated from words of $B$.
+Let $s\in B$ have the property that for all $r,t\in B$, either rstrst is not low, or rstrst contains a factor from $S$.
+Then a suffix of $\mathbf{w}$} is concatenated from words of $B-\{s\}$.
 5 Proof of the structure theorem
 We prove the structure theorem via an intricate case analysis.
-Each case is determined by a set $S\subseteq\Sigma_{3}^{*}$ where we assume that $S\subseteq F$, the set of factors omitted by $\mathbf{w}$}.
+Each case is determined by a set $S\subseteq F$, where $F$ is the set of ternary words which are not factors of $\mathbf{w}$}.
 As remarked in the previous section, it suffices to resolve the subcases $S=\{01,12,20,00\}$ and $S=\{11,22\}$.
-In a given case, labelled by a set $S$, we proceed in one of three ways. We either:
+In a given case, labelled by a set $S$, we proceed in one of three ways. We either
   1. 1.
@@
 If at some point all subcases have been resolved (and Hercules beats the hydra) the structure theorem has been proved.
-The concrete recipe for carrying out 1, 2, and 3 above is as follows:
+Let a set $S$ of forbidden factors be given. Say that a word $w$ is good if it is low and has no factor in $S$. The concrete recipe for carrying out 1, 2, and 3 above is as follows.
   1. A.
@@
   2. B.
-    Perform a backtrack search, looking for a low word $w$ of length 250 with no factors in $S$; if no such word exists, the case cannot arise.
+    Perform a backtrack search, looking for a good word $w$ of length 250. Word $w$ is a hypothetical prefix of $\mathbf{w}$}. If no such word $w$ exists, the case cannot arise.
   3. C.
@@
     1. a.
-      If no such $n$ is found: We look for a morphism satisfying the conditions of the Morphism Lemma or the Dual Morphism Lemma.
-To do this, we find factors $b$ of $w$, with $|b|$ up to length 3, such that $b$ has at most 3 recurrent return words in $\mathbf{w}$}.
-We form a morphism from these 3 return words, and consider it and its conjugates.
-      If one of the considered morphisms satisfies the conditions of the Morphism Lemma or the Dual Morphism Lemma the case is resolved.
-    2. b.
       If we find $n$ such that $w$ has more than $2n+1$ factors of length $n$, then we classify each such factor $s$ as needed or unneeded.
@@
       1. i.
-        If there are more than $2n+1$ needed factors then the condition on the complexity of $\mathbf{w}$} is violated and the case cannot arise.
+        If there are more than $2n+1$ needed factors then no low word of length 250 can be a prefix of $\mathbf{w}$}, since the condition on the complexity of $\mathbf{w}$} is violated.
+The case where $\mathbf{w}$} omits factors in $S$ cannot arise.
       2. ii.
         Otherwise, for each unneeded factor $s$ we form a subcase, replacing $S$ by $S\cup\{s\}$.
+    2. b.
+      If no such $n$ is found, we look for a morphism satisfying the conditions of the morphism lemma or the dual morphism lemma. To do this
+      1. i.
+        Consider the good words $b$ with $|b|\leq 3$.
+      2. ii.
+        Suppose that for some positive integer $d\leq 200$, every good word of length $d+1$ has $b$ as a factor. This can be checked by a finite search.
+      3. iii.
+        Then consider the set $B=\{bu\in\Sigma_{3}^{*}:bub$ is good, $|u|\leq d$, and whenever $p$ is a prefix of bubbub such that $b$ is a suffix of $p$, then $p=b$ or p=bub}p=bub\}.
+By the first parsing lemma, $\mathbf{w}$} is concatenated from words of $B$.
+      4. iv.
+        We form a morphism using the blocks in B if either:
+        • Set $B$ contains exactly 3 $b$-blocks;
+        • Set $B$ contains exactly 4 $b$-blocks, including a $b$-block $s$ such that for all $r,t\in B$, word rstrst is not good.
+In this case we replace $B$ by $B-\{s\}$.
+By the second parsing lemma, a final segment of $\mathbf{w}$} is concatenated from the $b$-blocks of $B$.
+      We form a morphism from the three $b$-blocks of $B$, and consider this morphism and its conjugates.
+If one of the considered morphisms satisfies the conditions of the morphism lemma or the dual morphism lemma the case is resolved.
 Remark 2.
@@
 Proof of the structure theorem.
-A python implementation of the above scheme in SageMath 9.3 (on a Microsoft Surface with Intel(R) Core(TM) i5-1035G4 CPU at 1.10GHz) resolved the case $S=\{11,22\}$ in about 15 seconds, and the case $S=\{01,12,20,00\}$ in about 4 seconds.
+A python implementation of the above scheme in SageMath 9.3 (on a Microsoft Surface with Intel(R) Core(TM) i5-1035G4 CPU at 1.10GHz) resolved the case $S=\{11,22\}$ in about 36 seconds, and the case $S=\{01,12,20,00\}$ in about 9 seconds.
+The code is available at https://github.com/drjamescurrie-ai/RACE-3/tree/main.
 ∎
@@
 We give the reasoning that was output for the case $S=\{01,12,20,00\}$ in an appendix.
-Further, the implementation kept track of all morphisms $h$ used.
-It may be the case that for some of these morphisms, $h({\mathbf{G}})$ is an eligible word with critical exponent $\mu$.
+Further, the implementation kept track of all the morphisms $h$ used.
+It may be the case that for some of these morphisms, the word $h({\mathbf{G}})$ is eligible and has critical exponent $\mu$.
 6 Proofs of Lemmas
@@
 The critical exponent of $\mathbf{u}$} is irrational, and therefore is not realized by any of its finite factors.
-Thus for any $\epsilon_{2}>0$ and any positive integer $N$, word $\mathbf{u}$} has a factor $uuu^{\prime}$ such that $u^{\prime}$ is a prefix of $u$ with $|u^{\prime}|>N$ and
+Thus for all $\epsilon_{2}>0$ and any positive integer $N$, the word $\mathbf{u}$} has a factor $uuu^{\prime}$ such that $u^{\prime}$ is a prefix of $u$ with $|u^{\prime}|>N$ and
 $\frac{|uuu^{\prime}|}{|u|}>\nu-\epsilon_{2}.$ (2)
@@
 We see that $\mathbf{w}$} contains factors with exponent arbitrarily close to $\nu$. The result follows. ∎
-We use two preliminary lemmas in the proof of the Morphism Lemma (Lemma 4):
+Proof of Lemma 3.
+Let $\mathbf{s}$} be a suffix of $\mathbf{w}$}. It suffices to show that $\mathbf{s}$} contains letter 0.
+Suppose $\mathbf{s}$} contains only letters 1 and 2. It must contain factors 12 and 21; otherwise $\mathbf{s}$} would end with $1^{\omega}$ or $2^{\omega}$ and therefore contain $5/2$ powers.
+Suffix $\mathbf{s}$} must also contain factor 11 (and, symmetrically, factor 22): If not, assume $\mathbf{s}$} starts with 1, replacing $\mathbf{s}$} with one of its suffixes if necessary.
+Then $\mathbf{s}$} is concatenated from the $1$-blocks of $\mathbf{w}$}.
+Since 222 is a $3$ power, the only possible $1$-blocks of $\mathbf{w}$} are 12 and 122.
+The $1$-block 12 can only be used once in this concatenation, or else it appears in the context $2\underline{12}12$, which is a $5/2$ power.
+Then a suffix is $(122)^{\omega}$, which contains $5/2$ powers.
+This is a contradiction.
+Thus 11 must indeed occur.
+So far we have shown that $\mathbf{s}$} contains factors 11, 12, 21, and 22.
+Since $\mathbf{w}$} contains $2(2)+1=5$ factors of length 2, there is exactly one length-2 factor of $\mathbf{w}$} containing a 0.
+It follows that the only 0 in $\mathbf{w}$} is its first letter.
+We thus see that $\mathbf{w}$} contains no factor of the form $a0$ where $a\in\Sigma_{3}$.
+Consider the following set of 10 length-4 words:
+$S=\{1121,1122,1211,1212,1221,2112,2121,2122,2211,2212\}.$
+Let $s$ and $t$ be distinct words of $S$.
+Suppose $u$ is a low word over $\Sigma_{3}$ such that the set of factors of $u$ does not include any of 00, 10, 20, $s$, or $t$.
+For each of the $\binom{10}{2}$ possibilities for $s$ and $t$, a backtrack search shows that the length of $u$ is at most 39.
+It then follows that $\mathbf{w}$} contains 9 out of 10 words of $S$ as factors.
+In addition,the word $\mathbf{w}$} contains a length-4 factor starting with 0.
+However, the word $\mathbf{w}$} is eligible, and cannot contain $10$ length-4 factors.
+This is a contradiction.
+∎
+We use two preliminary lemmas in the proof of the morphism lemma (Lemma 4).
 Lemma 8.
 Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low.
-Let $a,b,c\in\Sigma_{3}^{+}$, and suppose that a final segment of $\mathbf{w}$} has the form $g({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$, where $g=[a,b,c]$.
-Suppose further that:
+Let $a,b,c\in\Sigma_{3}^{+}$, and suppose that a suffix of $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$, where $h=[a,b,c]$.
+Suppose further that
   1. 1.
-    Word $b$ is a prefix of $a$, which is a prefix of $c$.
+    The word $b$ is a prefix of $a$, which is a prefix of $c$.
   2. 2.
@@
   3. 3.
-    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
-Then $\mathbf{u}$} doesn’t contain factors $00$, $11$, 2121, or $22$.
+    The words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
+Then $\mathbf{u}$} doesn’t contain any of the factors $00$, $11$, 2121, and $22$.
 Proof.
-We treat the factors one at a time. For each factor we rule out 1-letter right extensions of that factor:
+We treat the factors one at a time. For each factor we rule out 1-letter right extensions of that factor.
 Factor $11$: If $11$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=110$, $v=111$, or $v=112$.
-In each case, $g(v)$ has the prefix bbbbbb which is a $3$ power.
-Since $\mathbf{w}$} doesn’t contain $3$ powers, $11$ doesn’t occur in $\mathbf{u}$}.
+In each case, $h(v)$ has the prefix bbbbbb which is a $3$ power.
+Since $\mathbf{w}$} doesn’t contain $3$ powers, the factor $11$ doesn’t occur in $\mathbf{u}$}.
 Factor $00$: If $00$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=000$, $v=001$, or $v=002$.
-In each case $g(v)$ has the prefix aabaab since $b$ is a prefix of $a$ and $c$.
-However, aabaab is an $r$ power, $r\geq 5/2$, since $b$ is a prefix of $a$ and $|b|\geq|c|/2\geq|a|/2$.
+In each case $h(v)$ has the prefix aabaab since $b$ is a prefix of $a$ and $c$.
+However, aabaab is an $r$ power with $r\geq 5/2$, since $b$ is a prefix of $a$ and $|b|\geq|c|/2\geq|a|/2$.
 Factor $21$: If $21$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=211$, $v=210$, or $v=212$.
-In each case, $g(v)$ has the prefix cbbcbb.
-This has the suffix sbbsbb, which is an $r$ power, $r\geq 5/2$.
+In each case, $h(v)$ has the prefix cbbcbb, since $b$ is a prefix of $a$ and $c$.
+This has the suffix sbbsbb, since $s$ is a suffix of $c$.
+But sbbsbb is an $r$ power with $r\geq 5/2$.
 Factor $22$: If $22$ is a factor of $\mathbf{u}$}, then so is a factor $v$, where $v=220$, $v=221$, or $v=222$.
-In each case, $g(v)$ has the prefix ccbccb, which is an $r$ power, $r\geq 5/2$, since $b$ is a prefix of $c$ and $|b|\geq|c|/2$.
+In each case, $h(v)$ has the prefix ccbccb, which is an $r$ power with $r\geq 5/2$, since $b$ is a prefix of $c$ and $|b|\geq|c|/2$.
 ∎
 It is convenient to work with $\gamma^{2}=[012,02,0102]$ rather than simply $\gamma$.
-If $g=[a,b,c]$, then $g\circ\gamma^{2}=[abc,ac,abac]$, and each image of a letter under $g\circ\gamma^{2}$ ends in $c$.
-This allows us to define $g^{\prime}=c(g\circ\gamma^{2})c^{-1}=[cab,ca,caba]$.
+If $h=[a,b,c]$, then $h\circ\gamma^{2}=[abc,ac,abac]$, and each image of a letter under $h\circ\gamma^{2}$ ends in $c$.
+This allows us to define $h^{\prime}=c(h\circ\gamma^{2})c^{-1}=[cab,ca,caba]$.
 Lemma 9.
 Suppose that ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ is low.
-Suppose that $\mathbf{w}$} has the form $g({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $g=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
+Suppose that $\mathbf{w}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that
   1. i.
-    Word $b$ is a prefix of $a$, which is a prefix of $c$.
+    The word $b$ is a prefix of $a$, which is a prefix of $c$.
   2. ii.
@@
   3. iii.
-    Words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
-Then a final segment of $\mathbf{w}$} has the form $g(\gamma^{2}({\mathbf{v}}))$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$; equivalently, a final segment of $\mathbf{w}$} has the form $g^{\prime}({\mathbf{v}}))$ some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$, where $g^{\prime}=c(g\circ\gamma^{2})c^{-1}=[cab,ca,caba]$.
-Letting $A=g^{\prime}(0)$, $B=g^{\prime}(1)$, and $C=g^{\prime}(2)$, we have
+    The words $b$ and $c$ have a common suffix $s$ such that $|s|\geq|b|/2$.
+Then a suffix of $\mathbf{w}$} has the form $h(\gamma^{2}({\mathbf{v}}))$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$; equivalently, a suffix of $\mathbf{w}$} has the form $h^{\prime}({\mathbf{v}}))$ some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$, where $h^{\prime}=c(h\circ\gamma^{2})c^{-1}=[cab,ca,caba]$.
+Letting $A=h^{\prime}(0)$, $B=h^{\prime}(1)$, and $C=h^{\prime}(2)$, we have
   1. 1.
-    Word $B$ is a prefix of $A$, which is a prefix of $C$.
+    The word $B$ is a prefix of $A$, which is a prefix of $C$.
   2. 2.
@@
   3. 3.
-    Words $B$ and $C$ have a common suffix $S$ such that $|S|\geq|B|/2$.
+    The words $B$ and $C$ have a common suffix $S$ such that $|S|\geq|B|/2$.
 Proof.
-Without loss of generality, replacing $\mathbf{w}$} by one of its final segments if necessary, suppose that $\mathbf{w}$} has the form $g({\mathbf{u}})$.
+Replacing $\mathbf{w}$} by one of its suffixes if necessary, suppose that $\mathbf{w}$} has the form $h({\mathbf{u}})$.
 By Lemma 8, words $00$ and $11$ are not factors of $\mathbf{u}$}.
-Therefore, if $2$ is not a factor of some final segment of $\mathbf{u}$}, then that final segment is $(01)^{\omega}$ or $(10)^{\omega}$.
+Therefore, if $2$ is not a factor of some suffix of $\mathbf{u}$}, then that suffix is $(01)^{\omega}$ or $(10)^{\omega}$.
 This is impossible since $\mathbf{w}$} is low.
-We conclude that every final segment of $\mathbf{u}$} contains a $2$.
-Parse a final segment of $\mathbf{u}$} into blocks starting with $2$, and containing a single $2$.
+We conclude that every suffix of $\mathbf{u}$} contains a $2$.
+Parse a suffix of $\mathbf{u}$} into blocks starting with $2$, and containing a single $2$.
 Words $21$ and $22$ are not factors of $\mathbf{u}$}; thus the blocks begin with $20$.
@@
 The block $20101$ is impossible; it would imply a block $z=2010101$ or $z=201012$.
-In both cases, $g(z)$ begins with cababacababa, since $a$ is a prefix of $c$.
-However, ababaababa is an $r$ power, $r\geq 5/2$.
-It follows that a final segment of $\mathbf{u}$} is concatenated from blocks $20$, $201$, and $2010$.
-Therefore, a final segment of $\mathbf{w}$} is concatenated from blocks $ca=A$, $cab=B$, and $caba=C$, and has the form $g^{\prime}({\mathbf{v}})$ for some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
-We see that property (1) certainly holds. Also,
+In both cases, $h(z)$ begins with cababacababa, since $a$ is a prefix of $c$.
+However, ababaababa is an $r$ power with $r\geq 5/2$.
+It follows that a suffix of $\mathbf{u}$} is concatenated from blocks $20$, $201$, and $2010$.
+Therefore, a suffix of $\mathbf{w}$} is concatenated from blocks $ca=A$, $cab=B$, and $caba=C$, and has the form $h^{\prime}({\mathbf{v}})$ for some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
+We see that property 1 certainly holds. Also,
 $|C|=|caba|\leq 2|ca|=2|B|$
-since $|c|\geq|b|$, establishing (2).
-Finally, $C$ and $B$ have the common suffix $S=sa$. However,
+since $|c|\geq|b|$, establishing 2.
+Finally, the words $C$ and $B$ have the common suffix $S=sa$. However,
 $2|s|+|a|\geq|b|+|a|\geq 2|b|\geq|c|.$
@@
 $|S|=|sa|\geq|c|-|s|=|ca|-|sa|=|B|-|S|,$
-so that $2|S|\geq|B|$, establishing (3). ∎
+so that $2|S|\geq|B|$, establishing 3. ∎
 Proof of morphism lemma.
-Define morphisms $g_{n}$ recursively by $g_{1}=g$, and $g_{n+1}=cg_{n}\circ\gamma^{2}c^{-1}$.
-By induction on the previous lemma, for each positive integer $n$, a final segment of $\mathbf{w}$} has the form $g_{n}({\mathbf{v}})$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
-Having a final segment of the form $g_{n+1}({\mathbf{v}})$ is equivalent to having a final segment of the form $g_{n}(\gamma^{2}({\mathbf{v}}))$.
-By induction on $n$, having a final segment of the form $g_{n}({\mathbf{v}})$ is equivalent to having a final segment of the form $g(\gamma^{2n}({\mathbf{v}}))$.
-Thus a final segment of $\mathbf{w}$} has the form $g(\gamma^{2n}({\mathbf{v}}))$.
+Define morphisms $h_{n}$ recursively by $h_{1}=h$, and $h_{n+1}=ch_{n}\circ\gamma^{2}c^{-1}$.
+By induction on the previous lemma, for each positive integer $n$, a suffix of $\mathbf{w}$} has the form $h_{n}({\mathbf{v}})$, some ${\mathbf{v}}\in\Sigma_{3}^{\omega}$.
+Having a suffix of the form $h_{n+1}({\mathbf{v}})$ is equivalent to having a suffix of the form $h_{n}(\gamma^{2}({\mathbf{v}}))$.
+By induction on $n$, having a suffix of the form $h_{n}({\mathbf{v}})$ is equivalent to having a suffix of the form $h(\gamma^{2n}({\mathbf{v}}))$.
+Thus a suffix of $\mathbf{w}$} has the form $h(\gamma^{2n}({\mathbf{v}}))$.
+It follows that whenever $g$ is a prefix of $\mathbf{G}$}, then $h(g)$ is a factor of $\mathbf{w}$}, and the lemma follows.
 ∎
 Proof of dual morphism lemma.
-Replacing $\mathbf{z}$} by a final segment if necessary, suppose that $\mathbf{z}$} has the form $h({\mathbf{u}})$.
-Let $S=\{z\in\Sigma_{3}^{*}:h(z)\text{ is a factor of }{\mathbf{z}}\}.$ Let $S^{R}=\{z^{R}:z\in S\}$.
-Then $S^{R}$ is an infinite factorial language.
-Using König’s Tree Lemma, choose a word ${\mathbf{z}}^{\prime}\in\Sigma_{3}^{\omega}$ such that every prefix of $\mathbf{z}$$^{\prime}$ is in $S^{R}$.
+Replacing $\mathbf{z}$} by a suffix if necessary, suppose that $\mathbf{z}$} has the form $h({\mathbf{u}})$.
+Let $T=\{z\in\Sigma_{3}^{*}:h(z)\text{ is a factor of }{\mathbf{z}}\}.$ Let $T^{R}=\{z^{R}:z\in S\}$.
+Then $T^{R}$ is an right factorial language.
+Using König’s lemma, choose a word ${\mathbf{z}}^{\prime}\in\Sigma_{3}^{\omega}$ such that every prefix of $\mathbf{z}$$^{\prime}$ is in $T^{R}$.
 Define $h^{R}=[h(0)^{R},h(1)^{R},h(2)^{R}]$.
-Let ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime}).$ Then ${\mathbf{w}}={\mathbf{Z}}$ and $g=h^{R}$ satisfy the conditions of the Morphism Lemma, so that for every $g\in{\mathbf{G}}$, $h^{R}(g)$ is a factor of ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime})$.
-Let $p$ be a prefix of $\mathbf{z}$$^{\prime}$ such that $h^{R}(g)$ is a factor of $h^{R}(p).$ Since every prefix of $\mathbf{z}$$^{\prime}$ is in $S^{R}$, write $p=z^{R}$ where $z\in S$.
+Let ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime}).$ Then ${\mathbf{w}}={\mathbf{Z}}$ and $h^{\prime}=h^{R}$ satisfy the conditions of the morphism lemma, so that for every $g\in{\mathbf{G}}$, word $h^{R}(g)$ is a factor of ${\mathbf{Z}}=h^{R}({\mathbf{z}}^{\prime})$.
+Let $p$ be a prefix of $\mathbf{z}$$^{\prime}$ such that $h^{R}(g)$ is a factor of $h^{R}(p).$ Since every prefix of $\mathbf{z}$$^{\prime}$ is in $T^{R}$, write $p=z^{R}$ where $z\in T$.
 Then $h^{R}(g)$ is a factor of $h^{R}(z^{R})$.
 Thus $h(g^{R})$ is a factor of $h(z)$.
-By the definition of $S$, $h(z)$ is a factor of $\mathbf{z}$}, so that $h(g^{R})$ is a factor of $\mathbf{z}$}.
+By the definition of $T$, $h(z)$ is a factor of $\mathbf{z}$}, so that $h(g^{R})$ is a factor of $\mathbf{z}$}.
+∎
+Proof of first parsing lemma.
+Let $\mathbf{w}$} be an eligible right-infinite word which omits the words of $S$.
+Every factor of $\mathbf{w}$} of length $d+1$ is low, and omits words of $S$.
+By the condition of the lemma it cannot omit $b$.
+Thus $b$ appears infinitely often in $\mathbf{w}$}, so that a suffix can be concatenated from the $b$-blocks of $\mathbf{w}$}.
+These blocks have the form $bu$ where
+  • The word bubbub is a factor of $\mathbf{w}$};
+  • Whenever $p$ is a prefix of bubbub such that $b$ is a suffix of $p$, then $p=b$ or $p=bub$.
+Such a word $u$ is a factor of $\mathbf{w}$} and is therefore low and omits words of $S$.
+If $|u|>d$ it must therefore have $b$ as a factor.
+In this case, write $u=vbz$.
+Then $p=bvb$ is a prefix of bubbub distinct from $b$ and bubbub.
+This is a contradiction.
+∎
+Proof of second parsing lemma.
+Write ${\mathbf{w}}=b_{1}b_{2}b_{3}b_{4}\cdots$ such that the $b_{i}\in B$.
+Then for $j\geq 2$, if $b_{j}=s$ let $b_{j-1}=r$ and $b_{j+1}=t$.
+Since rstrst is a factor of $\mathbf{w}$} it is low and cannot contain a factor from $S$.
+This is a contradiction.
+We conclude that the suffix $b_{2}b_{3}b_{4}\cdots$ of $\mathbf{w}$} is concatenated from word of $B-\{s\}$.
 ∎
@@
 Consider the morphism which is the reverse of $\gamma$, namely $\gamma^{R}=[10,2,20]$. Let $\mathbf{G}$$^{\prime}$ be its fixed point
-$(\gamma^{R})^{\infty}(2)=20102102010\cdots.$
-Then the factors of $\mathbf{G}$$^{\prime}$ are precisely the reverses of the factors of $\mathbf{G}$}. One can easily adjust the proof of the Morphism Lemma to prove
+${\mathbf{G}}^{\prime}=(\gamma^{R})^{\infty}(2)=20102102010\cdots.$
+(An anonymous referee points out that ${\mathbf{G}}^{\prime}=\pi({\mathbf{p}})$ where $\mathbf{p}$} is the sequence studied in [11] and $\pi$ is the permutation of $[2,0,1]$ of $\Sigma_{3}$.
+See the entry for sequence A287072 in the On-Line Encyclopedia of Integer Sequences.) Then the factors of $\mathbf{G}$$^{\prime}$ are precisely the reverses of the factors of $\mathbf{G}$}.
+One can easily adjust the proof of the morphism lemma to prove
 Lemma 10 (Alternate Dual Morphism Lemma).
-Suppose that ${\mathbf{z}}\in\{0,1,2\}^{\omega}$ is low.
-Suppose that a final segment of $\mathbf{z}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\{0,1,2\}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that:
+Suppose that ${\mathbf{z}}\in\Sigma_{3}^{\omega}$ is low.
+Suppose that a suffix of $\mathbf{z}$} has the form $h({\mathbf{u}})$, some ${\mathbf{u}}\in\Sigma_{3}^{\omega}$ where $h=[a,b,c]$, some $a,b,c\in\Sigma_{3}^{+}$ such that
   1. 1.
-    Word $b$ is a suffix of $a$, which is a suffix of $c$.
+    The word $b$ is a suffix of $a$, which is a suffix of $c$.
   2. 2.
@@
   3. 3.
-    Words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
+    The words $b$ and $c$ have a common prefix $p$ such that $|p|\geq|b|/2$.
 Then if $g$ is any factor of $\mathbf{G}$$^{\prime}$, then $h(g)$ is a factor of $\mathbf{z}$}.
@@
 Theorem 4 (Sharpened Structure Theorem).
-Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be low, and let $\mathbf{w}$} have factor complexity $2n+1$. One of the following holds:
-  • For some non-erasing morphism $h$ in Figure 1, $h(g)$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$};
-  • For some non-erasing morphism $h$ in Figure 2, $h(g)$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$$^{\prime}$.
+Let ${\mathbf{w}}\in\Sigma_{3}^{\omega}$ be eligible. One of the following holds.
+  • For some non-erasing morphism $h$ in Figure 1 and a permutation $\pi$ of $\Sigma_{3}$, the word $\pi(h(g))$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$};
+  • For some non-erasing morphism $h$ in Figure 2 and some permutation $\pi$ of $\Sigma_{3}$, the word $\pi(h(g))$ is a factor of $\mathbf{w}$} whenever $g$ is a factor of $\mathbf{G}$$^{\prime}$.
 $h(0)$ $h(1)$ $h(2)$
@@
 $021201212$ $0212012$ $021201212012$
-$0201212012$ $0201212$ $020121201212$
-$0201212$ $02012$ $0201212012$
-$201021010$ $2010210$ $201021010210$
-$0102121021$ $0102121$ $010212102121$
-$0102121$ $01021$ $0102121021$
-$102012020$ $1020120$ $102012020120$
 $21021102211021102$ $2102110221102$ $2102110221102110221102$
 Figure 1: Potential morphisms $h$ such that some eligible word contains $h(g)$ for each factor $g$ of $\mathbf{G}$}.
+$h(0)$ $h(1)$ $h(2)$
 $10212$ $0212$ $0210212$
@@
 $1212010$ $12010$ $1201212010$
-$1210120$ $10120$ $101210120$
-$202012021$ $2012021$ $201202012021$
-$21020$ $1020$ $1021020$
-$0210202101$ $0202101$ $020210202101$
-$0202101$ $02101$ $0210202101$
-$1012102$ $12102$ $121012102$
-$101021012$ $1021012$ $102101021012$
-$0101202$ $01202$ $0120101202$
-$0120101202$ $0101202$ $010120101202$
-$010120102$ $0120102$ $012010120102$
 $221102210$ $21102210$ $2110221102210$
 Figure 2: Potential morphisms $h$ such that some eligible word contains $h(g)$ for each factor $g$ of $\mathbf{G}$$^{\prime}$.
-We note in passing that the third morphism in Figure 1, namely
-$h=[021201212,0212012,021201212012],$
-discovered by the computer implementation of the proof, is a conjugate of
+The morphisms in Figures 1 and 2 were output by the computer implementation of our proof.
+In fact, the implementation gave 26 different morphisms, each produced by applying a permutation to one of the listed morphisms.
+Some further simplification is possible.
+For example, consider the third morphism in Figure 1, namely
+$h=[021201212,0212012,021201212012].$
+This is a conjugate of
 $h_{1}=[201212021,2012021,201212012021].$
@@
 $\sigma(h_{1})=[012020102,0120102,012020120102]=\gamma^{4}.$
-Morphism $h$ could thus be replaced in the figure/proof by the identity morphism.
-However, we did not complicate our proof implementation by searching for simplifications and/or duplications of morphisms.
-The sharpened theorem says that the factor set of $\mathbf{w}$} is the same as the factor set of $h({\mathbf{G}})$ for a morphism in Figure 1, or $h({\mathbf{G}}^{\prime})$ for a morphism in Figure 2.
+Morphism $h$ could thus be replaced in Figure 1 by the identity morphism. We did not complicate our proof implementation by searching for simplifications/duplications of morphisms.
+Open Problem 1.
+Give a version of the sharpened structure theorem with an ‘aesthetically optimal’ list of morphisms.
+The sharpened theorem says that the set of factors of $\mathbf{w}$} is the same as the set of factors of $h({\mathbf{G}})$ for a morphism in Figure 1, or $h({\mathbf{G}}^{\prime})$ for a morphism in Figure 2.
 It is not known for which, if any, of the $h$ the words $h({\mathbf{G}})$ and $h({\mathbf{G}}^{\prime})$ have factor complexity $2n+1$, and what the critical exponents of these words are.
@@
 For each morphism $h$ in Figure 2, the word $h({\mathbf{G}}^{\prime})$ has factor complexity $2n+1$ and critical exponent $\mu$.
-Appendix 1: ‘Human readable’ proof of the resolution of the case $S=\{01,12,20,00\}$ of the Structure Theorem
-Our Python implementation outputs the following proof:
-We assume $F$ includes
-$\{01,12,20,00\}$
+Taken together with Carpi and De Luca’s result [5] for Sturmian words, we have the beginnings of a Dejean’s Theorem for Arnoux-Rauzy words; we know the minimal critical exponent in the cases $k=2$ and $k=3$.
+Open Problem 2.
+For each integer $k\geq 4$, give an Arnoux-Rauzy word over $\Sigma_{k}$ with the least critical exponent.
+Appendix: ‘Human readable’ proof of the resolution of the case $S=\{01,12,20,00\}$ of the structure theorem
+The Python implementation outputs the following proof:
+We assume $F$ includes [01, 12, 20, 00].
 Every good word longer than 25 must include factor(s)
-$\{0221,1021,1022,1102,2102,2110\}$
+[0221, 1021, 1022, 1102, 2102, 2110].
 Author’s comment: In step B of the algorithm, we found a length-250 low word $w$ omitting $\{01,12,20,00\}$.
@@
 Of these, 6 (announced above) are found to be needed factors in step C.b of the algorithm.
-Word w must omit a factor from
-$\{0210,0211,2210,2211\}$
+The word w must omit a factor from
+[0210, 0211, 2210, 2211].
 Author’s comment: These are the 4 unneeded factors found in step C.b of the algorithm.
@@
 This gives rise to 4 cases.
+Case 1: $F$ includes [01, 12, 20, 00, 0210]
+Case 2: $F$ includes [01, 12, 20, 00, 0211]
+Case 3: $F$ includes [01, 12, 20, 00, 2210]
+Case 4: $F$ includes [01, 12, 20, 00, 2211]
 Author’s comment: This is the case division of step C.b.ii.
-Case 1: $F$ includes
-$\{01,12,20,00,0210\}$
+Case 1: $F$ includes [01, 12, 20, 00, 0210].
 Every good word longer than 68 must include factor(s)
-$\{02110,02210,02211,10211,10221,11022,21021,21102,22102,22110\}$
-Word w must omit a factor from
-$\{11021,21022\}$
+[02110, 02210, 02211, 10211, 10221, 11022, 21021, 21102, 22102, 22110].
+The word w must omit a factor from
+[11021, 21022].
 This gives rise to 2 cases.
-  Case 1.1: $F$ includes
-$\{01,12,20,00,0210,11021\}$
+Case 1.1: $F$ includes [01, 12, 20, 00, 0210, 11021]
+Case 1.2: $F$ includes [01, 12, 20, 00, 0210, 21022]
+Case 1.1: $F$ includes [01, 12, 20, 00, 0210, 11021].
 Every good word longer than 81 must include factor(s)
-$\{02110221022,02110221102,02210211022,02210221102,02211022102,$
-10211022102,10211022110,10221021102,10221022110,10221102210,10211022102,10211022110,10221021102,10221022110,10221102210,
-11022102110,11022102211,11022110221,21021102210,21021102211,11022102110,11022102211,11022110221,21021102210,21021102211,
-21022110221,21102210211,21102210221,21102211022,21022110221,21102210211,21102210221,21102211022,
-22102110221,22102211022,22110221021,22110221022}22102110221,22102211022,22110221021,22110221022\}
-In this case w must omit factor 02110221021
+[02110221022, 02110221102, 02210211022, 02210221102, 02211022102, 10211022102, 10211022110, 10221021102, 10221022110, 10221102210, 11022102110, 11022102211, 11022110221, 21021102210, 21021102211, 21022110221, 21102210211, 21102210221, 21102211022, 22102110221, 22102211022, 22110221021, 22110221022].
+In this case w must omit the factor 02110221021.
 Author’s comment: Here, because only 1 unneeded factor was found in Step C.b, we don’t make a new subcase.
+The $210$-blocks are among [2102, 21021102, 210221102, 2102110221102].
 Author’s comment: It turns out that when we find a length-250 low word $w$ omitting the current list of missing factors, the complexity is $2n+1$ for $n=1,2,\ldots,20$.
-We therefore (in step C.a) search for a morphism, considering factors $b$ of $w$ of length up to 3, and finding the return words of $b$.
-The return words of 210 are among
-$\{2102,21021102,210221102,2102110221102\}$
-Word 2102 cannot occur twice in a concatenation of these
-Author’s comment: If 2102 occurs twice, it occurs in the context $s2102t$, where $s,t$ are two of the return words. However, one checks that each such word $s2102t$ contains an $r$ power, $r\geq 5/2$.
-Thus a final segment of w is concatenated from
-$\{21021102,210221102,2102110221102\}$
-Taking conjugates, a final segment is the image under morphism
+We therefore, in step C.a, search for a morphism, considering factors $b$ of $w$ with $|b|\leq 3$ and finding the $b$-blocks of $w$.
+The word 2102 cannot occur twice in a concatenation of these
+Author’s comment: If 2102 occurs twice, it occurs in the context $s2102t$, where $s,t$ are two of the $210$-blocks.
+However, one checks that each such word $s2102t$ contains an $r$ power with $r\geq 5/2$.
+Thus a suffix of w is concatenated from [21021102, 210221102, 2102110221102].
+Taking conjugates, a suffix is the image under morphism
+[221102210, 21102210, 2110221102210].
+which satisfies the dual morphism lemma.
+(This concludes Case 1.1).
+Author’s comment: Since $\mathbf{w}$} has a suffix of the form $h({\mathbf{u}})$, the dual morphism lemma shows that $\mathbf{w}$} has critical exponent at least $\mu$, and the current subcase is resolved.
+Case 1.2: $F$ includes [01, 12, 20, 00, 0210, 21022].
+The 210-blocks are
+[21021102, 2102110221102, 21021102211021102, 2102110221102110221102].
+The word 21021102 cannot occur twice in a concatenation of these.
+Thus a suffix of w is concatenated from
 [2102110221102, 21021102211021102, 2102110221102110221102]
-which satisfies the Dual Morphism Lemma.
-Author’s comment: Since $\mathbf{w}$} has a final segment of the form $h({\mathbf{u}})$, the Dual Morphism Lemma shows that $\mathbf{w}$} has critical exponent at least $\mu$, and the current subcase is resolved.
-Case 1.2: $F$ includes
-$\{01,12,20,00,0210,21022\}$
-The return words of 210 are:
-$\{21021102,2102110221102,21021102211021102,2102110221102110221102\}$
-Word 21021102 cannot occur twice in a concatenation of these.
-    Thus a final segment of w is concatenated from:
-$\{2102110221102,21021102211021102,2102110221102110221102\}$
-Taking conjugates, a final segment is the image under morphism
-[21021102211021102,2102110221102,2102110221102110221102][21021102211021102,2102110221102,2102110221102110221102]
+Taking conjugates, a suffix is the image under morphism
+[21021102211021102, 2102110221102, 2102110221102110221102].
 which satisfies the morphism lemma.
-    Case 2: $F$ includes
-$\{01,12,20,00,0211\}$
+(This concludes Case 1.2).
+(This concludes Case 1).
+Case 2: $F$ includes [01, 12, 20, 00, 0211].
 Every good word longer than 54 must include factor(s)
-$\{02102,02210,02211,10210,10221,11021,11022,21021,21022,21102,22102,22110\}$
-Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
+[02102, 02210, 02211, 10210, 10221, 11021, 11022, 21021, 21022, 21102, 22102, 22110].
+The word w has $12>2\times 5+1$ factors of length 5. This is impossible.
+(This concludes Case 2).
 Author’s comment: We found too many needed factors in step C.b.
-Case 3: $F$ includes
-$\{01,12,20,00,2210\}$
+Case 3: $F$ includes [01, 12, 20, 00, 2210].
 Every good word longer than 56 must include factor(s)
-$\{02102,02110,02211,10210,10211,10221,11021,11022,21021,21022,21102,22110\}$
-Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
-  Case 4: $F$ includes
-$\{01,12,20,00,2211\}$
+[02102, 02110, 02211, 10210, 10211, 10221, 11021, 11022, 21021, 21022, 21102, 22110].
+The word w has $12>2\times 5+1$ factors of length 5. This is impossible.
+(This concludes Case 3).
+Case 4: $F$ includes [01, 12, 20, 00, 2211].
 Every good word longer than 41 must include factor(s)
-$\{02102,02110,02210,10210,10211,10221,11021,11022,21021,21022,21102,22102\}$
-Word w has $12>2\times 5+1$ factors of length 5 . This is impossible.
+[02102, 02110, 02210, 10210, 10211, 10221, 11021, 11022, 21021, 21022, 21102, 22102].
+The word w has $12>2\times 5+1$ factors of length 5. This is impossible.
+(This concludes Case 4).
 Author’s comment: The proof in the present case is much shorter than for the resolution of $S=\{11,22\}$.
-In that longer resolution the computer frequently invokes step A of the algorithm (not used in the present case) avoid repetition of case analysis.
-Appendix 2: Python implementation of the case resolution software
-The following Python worksheet performs all the calculations mentioned in the paper:
-import math
-import time
-def complexityBreak(w):
-# Let the set of factors of w of length n be c[n]. If for some
-# 0<i<depth+1 we have |c[i]|>2i+1, return c[i] for the least such
-# i. Otherwise return 0.
-    depth=20
-    for i in range(1,depth+1):
-        c=[]
-        for j in range(len(w)-i+1):
-            if not(w[j:j+i]in c):
-                c.append(w[j:j+i])
-        if (len(c)>2*i+1):
-            c.sort()
-            return(c) #
-    return(0)
-def goodMorphism(L):
-# Does L=[a,b,c] satisfy the conditions of the Morphism Lemma?
-    def prefix(a,b):
-    # Does word b have prefix a?
-        if (len(b)<len(a)):
-            return(False)
-        else:
-            return(b[:len(a)]==a)
-    def goodsuff(a,c):
-    # Do c and a have a common suffix of length |a|/2?
-        for i in range(math.ceil(len(a)/2),len(a)+1):
-            if (a[-i:]==c[-i:]):
-                return(True)
-        return(False)
-    a=L[0]
-    b=L[1]
-    c=L[2]
-    return(prefix(a,b)and prefix(b,c)and goodsuff(a,c)
-           and (2*len(a)>=len(c)))
-def dualMorphism(L):
-# Does L=[a,b,c] satisfy the conditions of the Dual Morphism Lemma?
-    A=L[0][::-1]
-    B=L[1][::-1]
-    C=L[2][::-1]
-    L=[A,B,C]
-    return(goodMorphism(L))
-def morph(L):
-# Returns a conjugate of morphism L which satisfies the Morphism Lemma
-# or the Dual Morphism Lemma. If no such conjugate exists, returns 0.
-    def morphism_conjugates(L):
-    # Given a morphism L (a list of length 3) such that L[0]<=L[1]<=L[2],
-    # returns a list of the conjugates of L.
-        conjugates=[]
-        conjugates.append(L)
-        M=L
-    # To begin, M is a copy of L. We iteratively conjugate by (prefix)
-    # letters
-        i=0
-        while(i<len(L[0])):
-            if ((M[0][0]==M[1][0])and(M[1][0]==M[2][0])):
-                # M = [xa,xb,xc], so B=[ax,bx,cx] is a conjugate.
-                B=[]
-                for j in range(3):
-                    B.append(M[j][1:]+M[j][0])
-                M=B
-                conjugates.append(B)
-            i=i+1
-    # Again, M is a copy of L. We iteratively conjugate by (suffix)
-    # letters
-        M=L
-        i=1
-        while(i<len(L[0])):
-            if ((M[0][-1]==M[1][-1])and(M[1][-1]==M[2][-1])):
-                # M = [ax,bx,cx], so B=[xa,xb,xc] is a conjugate.
-                B=[]
-                for j in range(3):
-                    B.append(M[j][-1]+M[j][:-1])
-                M=B
-                conjugates.append(B)
-            i=i+1
-        return(conjugates)
-    C=morphism_conjugates(L)
-    for M in C:
-        if goodMorphism(M):
-            N=[M[1],M[0],M[2]]
-            return([1,N])
-        if dualMorphism(M):
-            N=[M[1],M[0],M[2]]
-            return([2,N])
-    return([0,[]])
-def fhpf(w): # Word w has no suffix of exponent 5/2 or greater
-    p=1 #potential period of suffix of exponent 5/2 or greater
-    while (5*p<=2*len(w)):
-        if (w[-3*p//2:]==w[-5*p//2:-p]):
-            return(False)
-        p=p+1
-    return(True)
-def test(w,forbidden_factors):
-# This returns true if no suffix of w is in forbidden_factors or has
-# exponent at least 5/2
-    for f in forbidden_factors:
-        if (len(w)>=len(f)):
-            if (w[-len(f):]==f):
-                return(False)
-    return(fhpf(w))
-def backtrack(target,forbidden_factors):
-# Call a ternary word _low if it contains no factor having exponent
-# 5/2 or greater. This routine returns a low word w starting
-# with 0 of length target such that w contains no factor in F. If no
-# such word exists, it returns the length of the longest low word
-# starting with 0 with no factor in forbidden_factors.
-    largest_letter=’2’
-    w=’0’
-    mx=1 # Greatest length attained so far
-    while(1==1):
-        success=test(w,forbidden_factors)
-        if (success):
-            if (len(w)>mx):
-                mx=len(w)
-            if (len(w)==target):
-                return(w)
-            else:
-                w=w+’0’
-        else:
-            while((len(w)>1) and w[-1]==largest_letter):
-                w=w[:-1]
-            if (len(w)==1):
-#               Search fails. Longest word has length mx.
-                return(mx)
-            else:
-                w=w[:-1]+chr(ord(w[-1])+1)
-def good(w,forbidden_factors):
-# Call a word _good if it is low and has no factor in
-# forbidden_factors.
-    for i in range(1,len(w)+1):
-        if not(test((w[:i]),forbidden_factors)):
-            return(False)
-    return(True)
-def doubler(n,forbidden_factors):
-# This returns the set of all good words of length n. It produces
-# the words recursively, testing concatenations of good words of
-# lengths ceil(n/2) and floor(n/2).
-    if (n==0):
-        return([’’])
-    if (n==1):
-        return([’0’,’1’,’2’])
-    G=[]
-    for s in doubler(math.ceil(n/2),forbidden_factors):
-        for t in doubler(n//2,forbidden_factors):
-            u=s+t
-            if good(s+t,forbidden_factors):
-                G.append(s+t)
-    return(G)
-def returns(b,depth,forbidden_factors):
-# Finds certain return words of b: words bu such that bub is
-# good, |u|<=depth, and there are exactly two instances of b in bub,
-# namely as prefix and suffix.
-    myReturns=[]
-    for i in range(depth+1):
-        forbidden_factors.append(b) # We seek good words u not
-                                    # containing b
-        U=doubler(i,forbidden_factors)
-        forbidden_factors.pop()
-        for u in U:
-            if good(b+u+b,forbidden_factors):
-                if ((b+u+b)[1:].find(b)==(len(b)+len(u)-1)):
-                    myReturns.append(b+u)
-    return(myReturns)
-def factors(w,n):
-# Returns the set of factors of w of length n
-    f=[]
-    for j in range(len(w)-n+1):
-        if not(w[j:j+n] in f):
-            f.append(w[j:j+n])
-    f.sort()
-    return(f)
-Morphisms=[]
-Dual_Morphisms=[]
-longest_b=1
-def findMorph(w,k,forbidden_factors):
-# Given word w and set of excluded words forbidden_factors, we
-# consider factors b # of w of length up to k. Suppose that every
-# good word of length v+1 contains b. We find the return words bu.
-# If there are exactly three return words (or four return words
-# of which the first can only occur once) then any good word has
-# a final segment concatenated from the three return words. We
-# form a morphism from these return words. We then test whether
-# one of its conjugates satisfies the Morphism Lemma or the Dual
-# Morphism Lemma,
-    global Morphisms
-    global Dual_Morphisms
-    global longest_b
-    n=1
-    while(n<=k):
-        Fact=factors(w,n)
-        for b in Fact:
-            forbidden_factors.append(b)
-            depth=backtrack(200,forbidden_factors)
-            forbidden_factors.pop()
-            if(isinstance(depth,int)):
-                # Longest good word not containing b has length
-                # depth
-                Returns=returns(b,depth,forbidden_factors)
-                if(len(Returns)>=3):
-                    if(len(Returns)<=4):
-                            myMorph=[Returns[-3],Returns[-2],Returns[-1]]
-                            N=morph(myMorph)
-                            if(N[0]==1):
-                            # This conjugate of myMorph satisfies
-                            # the Morphism Lemma. We check that
-                            # Returns[0] cannot appear twice
-                            # in a good concatenation of the return
-                            # words
-                                check=True
-                                for p in Returns:
-                                    for s in Returns:
-                                        if (good(p+Returns[0]+s,
-                                                 forbidden_factors)):
-                                            check=False
-                                if (check):
-                                    print(indent+’The return words ’+
-                                          ’of ’,b,’ are:’+’\\’+’\\’)
-                                    print(Returns,’\\’+’\\’)
-                                    print(indent+’Word ’,Returns[0],
-                                    ’cannot occur twice in a ’+
-                                    ’concatenation of these.\\’+’\\’)
-                                    print(indent+’Thus a final ’+
-                                    ’segment of w is concatenated ’+
-                                    ’from:’,’\\’+’\\’)
-                                    print(myMorph,’\\’+’\\’)
-                                    print(indent+’Taking conjugates,’+
-                                    ’ a final segment is the image ’+
-                                    ’under morphism’,’\\’+’\\’)
-                                    print(N[1],’\\’+’\\’)
-                                    if (not(N[1]in Morphisms)):
-                                        Morphisms.append(N[1])
-                                    print(indent+’which satisfies’+
-                                    ’ the Morphism Lemma.\\’+’\\’)
-                                    return()
-                            if(N[0]==2): # This conjugate of myMorph
-                                         # satisfies the Dual Morphism Lemma
-                                check=True
-                                for p in Returns:
-                                    for s in Returns:
-                                        if (good(p+Returns[0]+s,
-                                                 forbidden_factors)):
-                                            check=False
-                                if (check):
-                                    print(indent+’The return words’+
-                                    ’ of ’,b,’ are among ’,
-                                    Returns,’\\’+’\\’)
-                                    print(indent+’Word ’,Returns[0],
-                                    ’cannot occur twice in a ’+
-                                    ’concatenation of these’,’\\’+
-                                    print(indent+’Thus a final ’
-                                    +’segment of w is ’+
-                                    ’ concatenated from ’,myMorph,
-                                    print(indent+’Taking conjugates,’+
-                                    ’ a final segment is the image ’+
-                                    ’under morphism’,’\\’+’\\’)
-                                    print(N[1],’\\’+’\\’)
-                                    if (not(N[1]in Dual_Morphisms)):
-                                        Dual_Morphisms.append(N[1])
-                                    print(indent+’which satisfies’+
-                                    ’ the Dual Morphism ’+
-                                    ’Lemma.’,’\\’+’\\’)
-                                    return()
-        n=n+1
-        longest_b=max(longest_b,n)
-    print(indent+’No morphism found up to length ’, k,’\\’+’\\’)
-    return()
-indent = ’’
-def resolveCase(forbidden_factors,caseString):
-    global indent
-    global resolved_cases
-    global resolved_case_labels
-    indent=’~’
-    for i in range(len(caseString)):
-        indent=indent+’~’
-    k=3 # Maximum length of b considered in return words bu
-    if (caseString==’’):
-        print(indent+’We assume F includes ’,forbidden_factors,
-    w=backtrack(250,forbidden_factors)
-    if (isinstance(w,int)):
-        print(’No good word longer than ’,w,’ avoids these factors.’
-        F=[]
-        for f in forbidden_factors:
-            F.append(f)
-        resolved_cases.append(F)
-        resolved_case_labels.append(caseString)
-        print(resolved_cases)
-        print(resolved_case_labels)
-        return()
-#   The case being resolved is labelled by forbidden_factors.
-#   However, we may find that additional factors are also necessarily
-#   avoided in this case. We will sharpen our further backtracking
-# searches within this case by avoiding these additional factors.
-#   To put it another way, when there is exactly one unneeded factor,
-#   we would get one new subcase. Since there is no branching, it is
-#   more intuitive to simply say the subcase is the "same" case, but
-#   we specify additional avoided factors in factorsToAvoid. The "base
-#   case" is still labelled by the set forbidden_factors
-    factorsToAvoid=[]
-    for f in forbidden_factors:
-        factorsToAvoid.append(f)
-    unneeded_factors=[]
-    while(len(unneeded_factors)<2):
-        w=backtrack(250,factorsToAvoid)
-        N=complexityBreak(w)
-        if (N==0): #complexity is good
-            M=findMorph(w,k,factorsToAvoid)
-            F=[]
-            for f in forbidden_factors:
-                F.append(f)
-            resolved_cases.append(F)
-            resolved_case_labels.append(caseString)
-            return()
-        else:
-            Mx=0
-            needed_factors=[]
-            unneeded_factors=[]
-            for s in N:
-                factorsToAvoid.append(s)
-                w=backtrack(250,factorsToAvoid)
-                factorsToAvoid.pop()
-                if (isinstance(w,int)):
-                    Mx=max(Mx,w)
-                    needed_factors.append(s)
-                else:
-                    unneeded_factors.append(s)
-            if(len(needed_factors)>0):
-                print(indent+’Every good word longer than ’,Mx,
-                ’must include factor(s)’,’\\’+’\\’)
-                print(needed_factors,’\\’+’\\’)
-                if(len(needed_factors)>2*len(N[0])+1):
-                    print(indent+’Word w has ’,len(needed_factors),
-                    ’> 2 x’,len(N[0]),’+ 1 factors of length ’,
-                    len(N[0]) ,’. This is impossible.’,’\\’+’\\’)
-                    F=[]
-                    for f in forbidden_factors:
-                        F.append(f)
-                    resolved_cases.append(F)
-                    resolved_case_labels.append(caseString)
-                    return()
-            if (len(unneeded_factors)==1):
-                print(indent+’In this case w must omit factor ’,
-                unneeded_factors[0],
-                s=unneeded_factors[0]
-                factorsToAvoid.append(s)
-    unneeded_factors.sort()
-    j=0
-    print(indent+’Word w must omit a factor from’,’\\’+’\\’)
-    print(unneeded_factors,’\\’+’\\’)
-    print(indent+’This gives rise to ’,len(unneeded_factors),
-    ’ cases.’,’\\’+’\\’)
-    for s in unneeded_factors:
-        if (caseString==’’):
-            subcase=chr(ord(’1’)+j)
-        else:
-            subcase=caseString+’.’+chr(j+49)
-        j=j+1
-        factorsToAvoid.append(s)
-        print(indent+’Case ’+subcase+’: F includes ’,factorsToAvoid,
-        # Test whether case was resolved previously
-        resolved=False
-        for c in resolved_cases:
-            resolved=True
-            for d in c:
-                if not(d in factorsToAvoid):
-                    resolved=False
-                    break
-            if (resolved):
-                i=resolved_cases.index(c)
-                print(’This was previously resolved in Case ’+
-                resolved_case_labels[i]+’.’,’\\’+’\\’)
-                break
-        if(not(resolved)):
-            resolveCase(factorsToAvoid,subcase)
-            factorsToAvoid.pop()
-    F=[]
-    for f in forbidden_factors:
-        F.append(f)
-    resolved_cases.append(F)
-    resolved_case_labels.append(caseString)
-    return()
-print(’Here is the resolution of S={22,11}:’)
-print(’ ’)
-resolved_cases=[]
-resolved_case_labels=[]
-startTime=time.time()
-resolveCase([’22’,’11’],’’)
-endTime=time.time()
-print(’Computation took ’,(endTime-startTime),’ seconds.’)
-print(’ ’)
-print(’ ’)
-print(’ ’)
-print(’Here is the resolution of S={01,12,20,00}:’)
-print(’ ’)
-resolved_cases=[]
-resolved_case_labels=[]
-startTime=time.time()
-resolveCase([’01’,’12’,’20’,’$0$’],’’)
-endTime=time.time()
-print(’Computation took ’,(endTime-startTime),’ seconds.’)
-print(’ ’)
-print(’ ’)
-print(’ ’)
-print(’Here are the morphisms satisfying the Morphism Lemma used ’+
-’in the case resolutions:’)
-print(’ ’)
-for m in Morphisms:
-    print(m)
-print(’ ’)
-print(’ ’)
-print(’ ’)
-print(’Here are the morphisms satisfying the Dual Morphism Lemma ’+
-’used in the case resolutions:’)
-print(’ ’)
-for m in Dual_Morphisms:
-    print(m)
-print(’ ’)
-print(’ ’)
-print(’ ’)
-print(’Here is the length of the longest word in the backtracks of’+
-’ Lemma 2:’)
-S=[’1121’, ’1122’, ’1211’, ’1212’, ’1221’, ’2112’,
-’2121’, ’2122’, ’2211’, ’2212’]
-F=[’10’,’20’,’$0$’]
-longest=0
-for s in S:
-        for t in S:
-            if (s!=t):
-                F.append(s)
-                F.append(t)
-                w=backtrack(500,F)
-                longest=max(w,longest)
-                F.pop()
-                F.pop()
-print(’ ’)
-print(longest)
+In that longer resolution the computer frequently invokes step A of the algorithm (not used in the present case) to avoid repeating case analysis which was already carried out.